### v2

Abstract
The Kirkwood superposition is a well-known tool in statistical physics to approximate the $n$-point correlation functions for $n\geq 3$ in terms of the density $\rho$ and the radial distribution function $g$ of the underlying system. However, it is unclear whether these approximations are themselves the correlation functions of some point process. If they are, this process is called the Kirkwood closure process. For the case that $g$ is the negative exponential of some nonnegative and regular pair potential $u$ existence of the the Kirkwood closure process was proved by Ambartzumian and Sukiasian. This result was generalized to the case that $u$ is a locally stable and regular pair potential by Kuna, Lebowitz and Speer, provided that $\rho$ is sufficiently small. In this work, it is shown that it suffices for $u$ to be stable and regular to ensure the existence of the Kirkwood closure process. Furthermore, for locally stable $u$ it is proved that the Kirkwood closure process is Gibbs and that the kernel of the GNZ-equation satisfies a Kirkwood-Salsburg type equation.
keywords:
Realizability, Point processes, Gibbs point processes, Kirkwood-Salsburg equations, radial distribution function
{AMS}
82B21, 60G55
1 Introduction
In classical statistical physics, point processes are often used to describe the distribution of interacting particles in equilibrium. Often, so-called Gibbs measures are used. In these models the energy of a configuration of particles is calculated via some interaction potentials and a configuration is more likely to be observed when the associated energy is low. However, in general, it is not possible to measure these interaction potentials nor calculate them easily from given snapshots of these configurations, see e.g. [3].
In practice, the available data are the so-called $n$-point correlation functions $\rho^{(n)}$ of the underlying point process. However, while it is possible to calculate them for arbitrary $n$, these calculations get very computationally expensive as soon as $n>2$ as good statistics require long simulation times and $n$-tuples of particles have to be counted. Thus, commonly the Kirkwood superposition approximation, introduced by Kirkwood in [4], is used, cf. [3], to approximate the higher-order correlation functions, i.e.
$\rho^{(n)}(\boldsymbol{{x}}_{n})\approx\rho^{n}\prod_{1\leq i<j\leq n}g(x_{i}-x_{j}).$ (1)
Here $\boldsymbol{{x}}_{n}=(x_{1},\dots,x_{n})$ and $g=\rho^{(2)}/\rho^{2}$ is the so-called radial distribution function of the point process. In [1] the question has been raised whether there is a point process $\mathsf{K}$ whose correlation functions are given by the right-hand side of (1). This means the closed form expression of the correlation functions of the process $\mathsf{K}$ are given by the Kirkwood superposition, thus this point process $\mathsf{K}$ is called the Kirkwood closure process. In this work sufficient conditions for the existence of $\mathsf{K}$ are investigated.
This question is related to an interesting inverse problem, namely, a realizability problem for point processes, see [6]:
„Given $\rho>0$ and a nonnegative function $g$, does there exist a point process with density $\rho$ and radial distribution function $g$?“
The Kirkwood closure process is one possible ansatz for the solution of this problem.
For the case that $g\leq 1$ Ambartzumian and Sukiasian showed in [1] that the Kirkwood closure process exists when $\rho$ is small enough. Later, using a different technique this result was extended by Kuna, Lebowitz and Speer in [6].
In the language of statistical mechanics Ambartzumian and Sukiasian showed the existence of the Kirkwood closure when $g=e^{-u}$ where $u$ is some nonnegative and regular pair potential and Kuna, Lebowitz and Speer extended the result for the case that $u$ is a pair potential which is locally stable (e.g. when $u$ has a hard-core, i.e. $u=+\infty$ around the origin) and regular. In this work a connection between the well-known Kirkwood-Salsburg equations and the Kirkwood closure process is used to show existence of the latter when $u$ is a stable and regular pair potential. In fact, the so-called Janossy densities of the Kirkwood closure process are (up to a factor) the solutions of the Kirkwood-Salsburg equations for a negative activity. In particular, this solution has many well-known properties, cf. [9].
The outline is as follows: After introducing the setting in Section 2, the existence of the Kirkwood closure is proved in Section 3. In Section 4 the Gibbsianness of the Kirkwood closure is discussed and in the last Section generalizations of to higher order closures are discussed.
2 Setting
2.1 The Kirkwood closure process
Any probability measure $\mathsf{P}$ on the space of configurations
$\Gamma\,=\,\bigl{\{}\,\gamma\subset\mathbb{R}^{d}\ \Big{|}\ \Delta\subset\mathbb{R}^{d}\text{ bounded}\,\Rightarrow\,N_{\Delta}(\gamma)<+\infty\bigr{\}}\,,$
equipped with the $\sigma$-algebra $\mathcal{F}:=\sigma(N_{\Delta}\mid\Delta\subset\mathbb{R}^{d}\text{ bounded})$ is called a point process. Here $N_{\Delta}(\gamma)=\#(\gamma_{\Delta})$ ($\gamma_{\Delta}=\gamma\cap\Delta$) is the number of elements of $\gamma$ in $\Delta$. For some bounded set $\Lambda\subset\mathbb{R}^{d}$ the configurations in $\Lambda$ are denoted by $\Gamma_{\Lambda}=\{\gamma\in\Gamma\mid\gamma\subset\Lambda\}$ and $\Gamma_{0}=\{\gamma\in\Gamma\mid\#\gamma<+\infty\}$ denotes the space of finite configurations. The elements of a family of symmetric functions $(j^{(n)}_{\Lambda})_{n\geq 0,\,\Lambda\subset\mathbb{R}^{d}\text{ bounded}}$ are called the Janossy densities of $\mathsf{P}$, if for every $F\colon\Gamma\to[0,+\infty)$ there holds
$\int_{\Gamma}F(\gamma_{\Lambda})\mathop{}\!\mathrm{d}\mathsf{P}(\gamma)=\sum_{n=0}^{\infty}\frac{1}{n!}\int_{\Lambda^{n}}F(\{\boldsymbol{{x}}_{n}\})j_{\Lambda}^{(n)}(\boldsymbol{{x}}_{n})\mathop{}\!\mathrm{d}\boldsymbol{{x}}_{n}$ (2)
where the term for $n=0$ is understood to be $F(\emptyset)j_{\Lambda}^{(0)}$. If the Janossy densities of a point process exist, they are unique up to (Lebesgue) null-sets and determine $\mathsf{P}$ completely.
The elements of a family of symmetric functions $(\rho^{(n)})_{n\in\mathbb{N}}$ are called the correlation functions of $\mathsf{P}$, if for every $n\in\mathbb{N}$ and $F\colon(\mathbb{R}^{d})^{n}\to[0,+\infty)$ there holds
$\int_{\Gamma}\sum_{x_{1},\dots,x_{n}\in\gamma\atop x_{i}\neq x_{j}}F(\boldsymbol{{x}}_{n})\mathop{}\!\mathrm{d}\mathsf{P}(\gamma)=\int_{(\mathbb{R}^{d})^{n}}F(\boldsymbol{{x}}_{n})\rho^{(n)}(\boldsymbol{{x}}_{n})\mathop{}\!\mathrm{d}\boldsymbol{{x}}_{n}.$ (3)
Also note the formula
$\rho^{(n)}(\boldsymbol{{x}}_{n})=\sum_{k=0}^{\infty}\frac{1}{k!}\int_{\Lambda^{k}}j_{\Lambda}^{(n+k)}(\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k},\qquad\boldsymbol{{x}}_{n}\in\Lambda^{n}.$ (4)
Here $j_{\Lambda}^{(n+k)}(\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})=j_{\Lambda}^{(n+k)}(x_{1},\dots,x_{n},y_{1},\dots,y_{k})$ for brevity. If the point process $\mathsf{P}$ is stationary then the correlation functions are translationally invariant, and one can write $\rho^{(n)}=\rho^{n}g^{(n)}$ for appropriate functions $g^{(n)}$ depending on $n-1$ variables, where $\rho$ is the so-called intensity or density of the point process. For $n=2$ the function $g=g^{(2)}$ is the so-called radial distribution function.
As mentioned in the introduction, a point process $\mathsf{K}_{\varsigma,\phi}$ is called Kirkwood closure process, if it has correlation functions and there is a $\varsigma>0$ and an even nonnegative function $\phi\colon\mathbb{R}^{d}\to[0,+\infty)$ such that
$\rho^{(n)}(\boldsymbol{{x}}_{n})=\varsigma^{n}\prod_{1\leq i<j\leq n}\phi(x_{i}-x_{j}),$ (5)
where the empty product is understood to be equal to one. In particular, this means that for the Kirkwood closure the approximation (1) is an equality. The existence of the Kirkwood closure will be discussed in Section 3.
The correlation functions $(\rho^{(n)})_{n\geq 1}$ of a point process $\mathsf{P}$ satisfy Ruelle’s bound, if there is a $\xi>0$ such that
$\rho^{(n)}(\boldsymbol{{x}}_{n})\leq\xi^{n}.$ ($\mathcal{R}_{\xi}$)
In this case it is said that $\mathsf{P}$ satisfies condition ($\mathcal{R}_{\xi}$). Any point process $\mathsf{P}$ satisfying condition ($\mathcal{R}_{\xi}$) has a number of nice properties. Firstly, in this case the correlation functions determine $\mathsf{P}$ uniquely, see [5]. Secondly, $\mathsf{P}$ is supported on a set of „nice“ configurations. Namely, any point process $\mathsf{P}$ satisfying condition ($\mathcal{R}_{\xi}$) is supported on the tempered configurations
$\Gamma_{*}:=\bigcup_{M\geq 1}\bigcap_{n\geq 0}\left\{\gamma\in\Gamma\mid N_{\Delta_{n}}(\gamma)\leq M\lambda\!\!\!\hskip 0.5pt\lambda(\Delta_{n})\right\}$ (6)
where $\Delta_{n}=\{x\in\mathbb{R}^{d}\mid n\leq|x|<n+1\}$ and $\lambda\!\!\!\hskip 0.5pt\lambda(\cdot)$ denotes the Lebesgue measure on $\mathbb{R}^{d}$, i.e. $\mathsf{P}(\Gamma_{*})=1$, cf. [5]. In this case $\mathsf{P}$ is called tempered. Lastly, for any point process satisfying condition ($\mathcal{R}_{\xi}$) the inverse to (4) holds, i.e. for any bounded $\Lambda\subset\mathbb{R}^{d}$ there holds
$j_{\Lambda}^{(n)}(\boldsymbol{{x}}_{n})=\sum_{k=0}^{\infty}\frac{(-1)^{k}}{k!}\int_{\Lambda^{k}}\rho^{(n+k)}(\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k},\qquad\boldsymbol{{x}}_{n}\in\Lambda^{n}$ (7)
where for $n=0$ the term $\rho^{(0)}=1$, see e.g. [5]. In fact, (7) can also be used to define a point process:
Theorem A
[Lenard [7]] Let $(\rho^{(n)})_{n\geq 1}$ be a family of nonnegative symmetric functions that satisfy ($\mathcal{R}_{\xi}$) for some $\xi>0$ such that for all $n\in\mathbb{N}$, all bounded $\Lambda\subset\mathbb{R}^{d}$ and all $\boldsymbol{{x}}_{n}\in\Lambda^{n}$
$\sum_{k=0}^{\infty}\frac{(-1)^{k}}{k!}\int_{\Lambda^{k}}\rho^{(n+k)}(\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}\geq 0$ (8)
and
$1+\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k!}\int_{\Lambda^{k}}\rho^{(k)}(\boldsymbol{{y}}_{k})\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}\geq 0$ (9)
Then there exists a point process $\mathsf{P}$ with correlation functions $(\rho^{(n)})_{n\geq 1}$.
The conditions (8) and (9) are called Lenard positivity. In general, it is not easy to check whether a family $(\rho^{(n)})_{n\geq 1}$ satisfies the Lenard positivity condition. However, for the correlation functions of the Kirkwood closure process sufficient conditions for Lenard positivity have been given. First, by Ambartzumian and Sukiasian in [1] and later these were generalized by Kuna, Lebowitz and Speer in [6]. Ambartzumian and Sukiasian relied on an approach using a cluster expansion and Kuna et al. used an ansatz related to the so-called Mayer-Montroll equations, which are both well-known tools from classical statistical mechanics. As previously mentioned, in this work an approach using the Kirkwood Salsburg equations is used to extend their results.
2.2 The Kirkwood-Salsburg operator
The Kirkwood-Salsburg equations are a well-known tool for grand-canonical Gibbs measures, cf. [9]. Let $u\colon\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$ be an even function bounded from below to which a translationally invariant Hamiltonian $H\colon\Gamma_{0}\to\mathbb{R}\cup\{+\infty\}$ is associated by
$H(\gamma)=\frac{1}{2}\sum_{x\neq y\in\gamma}u(x-y).$ (10)
The function $u$ is called a translationally invariant pair potential. For $\beta>0$ the Mayer function of $u$ at inverse temperature $\beta$ is defined as
$f_{\beta}\,(x)=e^{-\beta u(x)}-1.$ (11)
Throughout it is assumed that $u$ is regular, i.e. that
$C_{\beta}\,(u):=\int_{\mathbb{R}^{d}}\left|f_{\beta}\,(x)\right|\mathop{}\!\mathrm{d}x<+\infty$ (12)
for all $\beta>0$. In fact, if there is a $\beta_{0}$ such that $C_{\beta_{0}}(u)<+\infty$, then $C_{\beta}\,(u)$ is finite for all $\beta>0$, cf. [9]. It will further be assumed that the pair potential $u$ (and thus the Hamiltonian $H$) is stable, meaning there is a $B>0$ such that
$H(\gamma)\geq-B\#\gamma.$ (13)
Remark 2.1.
A sufficient condition for $u$ to be stable and regular, is that $u$ is of Lennard-Jones type, i.e. that there exist $r_{0}>0$, $\alpha>d$, and $C>c>0$ such that
$u(x)\geq c|x|^{-\alpha},\quad|x|<r_{0},\quad\text{ and }\quad|u(x)|\leq C|x|^{-\alpha},\quad|x|\geq r_{0}.$
The interaction between $\eta\in\Gamma_{0}$ and $\gamma\in\Gamma$ is defined by
$W(\eta\mid\gamma):=\begin{dcases}\sum_{x\in\eta,y\in\gamma}u(x-y),\qquad\,\,\text{ if }\quad\sum_{x\in\eta,y\in\gamma}|u(x-y)|<+\infty\\ +\infty,\qquad\qquad\qquad\quad\text{ otherwise.}\end{dcases}$ (14)
For two finite configurations $\eta,\gamma\in\Gamma_{0}$, there holds
$H(\eta\cup\gamma)=H(\eta)+W(\eta\mid\gamma)+H(\gamma).$ (15)
Since $u$ is assumed to be a stable pair potential, every configuration $\boldsymbol{{x}}_{n}$ has an element $x_{i_{*}}$ with $i_{*}=i_{*}(\boldsymbol{{x}}_{n})\in\{1,\dots,n\}$ such that
$W(\{x_{i_{*}}\}\mid\{\boldsymbol{{x}}_{n-1}^{\prime}\})=\sum_{\begin{subarray}{c}i=1\\ i\neq i_{*}\end{subarray}}^{n}u(x_{i}-x_{i_{*}})\geq-2B$ (16)
where $\boldsymbol{{x}}_{n-1}^{\prime}=(x_{1},\dots,x_{i_{*}-1},x_{i_{*}+1},\dots,x_{n})$ is the configuration of the remaining elements, cf. [9]. In case there is more than one possible choice such that (16) holds, let $i_{*}$ be the smallest index with this property.
For $\zeta>0$, let
$E_{\zeta}:=\left\{\boldsymbol{\theta}=(\theta^{(n)})_{n\geq 1}\;\middle|\;\theta^{(n)}\colon(\mathbb{R}^{d})^{n}\to\mathbb{C},\,\,\,\|\boldsymbol{\theta}\|_{\zeta}<+\infty\right\}$ (17)
be the Banach space of sequences of complex $L^{\infty}$-functions with an increasing number of variables, for which the norm
$\|\boldsymbol{\theta}\|_{\zeta}:=\sup_{n\geq 1}\left(\zeta^{n}\|\theta^{(n)}\|_{\infty}\right)$
is finite and introduce the Kirkwood-Salsburg operator $\boldsymbol{K}\colon E_{C_{\beta}\,(u)}\to E_{e^{2\beta B}C_{\beta}\,(u)}$ as
$(\boldsymbol{K}\boldsymbol{\theta})^{(1)}(x)=\sum_{k=1}^{\infty}\frac{1}{k!}\int_{(\mathbb{R}^{d})^{k}}\prod_{i=1}^{k}f_{\beta}\,(x-y_{i})\theta^{(k)}(\boldsymbol{{y}}_{k})\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}$ (18)
and for $n\geq 1$ as
$(\boldsymbol{K}\boldsymbol{\theta})^{(n+1)}(x,\boldsymbol{{x}}_{n})=e^{-\beta W(\{x\}\mid\{\boldsymbol{{x}}_{n}\})}\left(\theta^{(n)}(\boldsymbol{{x}}_{n})+\sum_{k=1}^{\infty}\frac{1}{k!}\int_{(\mathbb{R}^{d})^{k}}\prod_{j=1}^{k}f_{\beta}\,(x-y_{j})\theta^{(n+k)}(\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}\right).$ (19)
Defining the permutation operator $\boldsymbol{\Pi}\colon E_{C_{\beta}\,(u)}\to E_{C_{\beta}\,(u)}$ by $(\boldsymbol{\Pi}\boldsymbol{\theta})^{(n)}(\boldsymbol{{x}}_{n})=\theta^{(n)}(x_{i_{*}},\boldsymbol{{x}}_{n-1}^{\prime})$, one finds by (16) that
$\sup_{n\geq 1}C_{\beta}\,(u)^{n}\|(\boldsymbol{\Pi}\boldsymbol{K}\boldsymbol{\theta})^{(n)}\|_{\infty}$ $\leq C_{\beta}\,(u)^{n}e^{2\beta B}\sum_{k=0}^{\infty}\frac{1}{k!}\int_{(\mathbb{R}^{d})^{k}}\prod_{j=1}^{k}|f_{\beta}\,(x_{i_{*}}-y_{j})|C_{\beta}\,(u)^{-n-k+1}\|\boldsymbol{\theta}\|\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}$
$\leq e^{2\beta B+1}C_{\beta}\,(u)$
and thus $\boldsymbol{\Pi}\boldsymbol{K}\colon E_{C_{\beta}\,(u)}\to E_{C_{\beta}\,(u)}$ is well-defined with $\|\boldsymbol{\Pi}\boldsymbol{K}\|_{E_{C_{\beta}\,(u)}\to E_{C_{\beta}\,(u)}}\leq e^{2\beta B+1}C_{\beta}\,(u)$. Lastly, for some bounded set $\Lambda\subset\mathbb{R}^{d}$ let $\boldsymbol{\chi}_{\Lambda}\colon E_{C_{\beta}\,(u)}\to E_{C_{\beta}\,(u)}$ be the projection operator
$\boldsymbol{\chi}_{\Lambda}\colon$ $E_{C_{\beta}\,(u)}\to E_{C_{\beta}\,(u)}$
$\boldsymbol{\theta}\mapsto\boldsymbol{\chi}_{\Lambda}\boldsymbol{\theta}=(\mathds{1}_{\Lambda^{n}}\theta^{(n)})_{n\geq 1},$
$\mathbf{I}\colon E_{C_{\beta}\,(u)}\to E_{C_{\beta}\,(u)}$ be the identity and $\boldsymbol{e}_{1}=(e_{1}^{(n)})_{n\geq 1}$ be the vector in $E_{C_{\beta}\,(u)}$ with $e^{(1)}_{1}\equiv 1$ and $e^{(n)}_{1}\equiv 0$ for $n\geq 2$.
For a given $z\in\mathbb{C}$ and bounded $\Lambda\subset\mathbb{R}^{d}$ consider the finite volume Kirkwood-Salsburg equations defined by
$(\mathbf{I}-z\boldsymbol{\chi}_{\Lambda}\boldsymbol{\Pi}\boldsymbol{K})\boldsymbol{\theta}=z\boldsymbol{\chi}_{\Lambda}\boldsymbol{e}_{1}.$ (20)
In the context of statistical mechanics $z$ (usually $z>0$) is called the activity of the grand-canonical ensemble associated to $(\beta,z,u)$. It is well-known that for $z\in B_{z_{0}}:=\{z\in\mathbb{C}\mid|z|<z_{0}\}$ where
$z_{0}:=\left(e^{2\beta B+1}C_{\beta}\,(u)\right)^{-1}$ (21)
there is a unique solution to (20) which can be developed into a Neumann-series, i.e. the solution is given by
$\boldsymbol{\theta}_{\Lambda}(z)=(\mathbf{I}-z\boldsymbol{\chi}_{\Lambda}\boldsymbol{\Pi}\boldsymbol{K})^{-1}z\boldsymbol{\chi}_{\Lambda}\boldsymbol{e}_{1}=\sum_{k=0}^{\infty}(z\boldsymbol{\chi}_{\Lambda}\boldsymbol{\Pi}\boldsymbol{K})^{k}z\boldsymbol{\chi}_{\Lambda}\boldsymbol{e}_{1}.$ (22)
In particular this means that for each $n$ and $x_{1},\dots,x_{n}\in\Lambda$ the function $\theta_{\Lambda}^{(n)}(z;\boldsymbol{{x}}_{n})$ is an analytic function on $B_{z_{0}}$. Furthermore, the solution of (20) can be written down explicitly using the grand canonical partition function
$\Xi_{\Lambda}(z)=1+\sum_{k=1}^{\infty}\frac{z^{k}}{k!}\int_{\Lambda^{k}}e^{-\beta H(\{\boldsymbol{{y}}_{k}\})}\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}.$ (23)
As shown by Ruelle, see [9], $\Xi_{\Lambda}(z)\neq 0$ for $z\in B_{z_{0}}$, which implies that
$\theta_{\Lambda}^{(n)}(z;\boldsymbol{{x}}_{n})=\frac{1}{\Xi_{\Lambda}(z)}\sum_{k=0}^{\infty}\frac{z^{n+k}}{k!}\int_{\Lambda^{k}}e^{-\beta H(\{\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k}\})}\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}.$ (24)
The argument that $\Xi_{\Lambda}(z)\neq 0$ by Ruelle is as follows: For $z>0$ and $n=1$ integration of (24) with respect to $x$ and differentiation of (23) with respect to $z$ shows that
$\int_{\Lambda}\theta^{(1)}_{\Lambda}(z;x)\mathop{}\!\mathrm{d}x=z\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}z}\log\Xi_{\Lambda}(z).$
Since by (22) the left-hand side is analytic in $B_{z_{0}}$ this implies that the right-hand can also be continued as an analytic function, meaning $\Xi_{\Lambda}(z)$ does not have any zeros in $B_{z_{0}}$. Using an argument in the same spirit in Section 3, it is shown that for fixed $x_{1},\dots,x_{n}\in\Lambda^{n}$ either $\theta^{(n)}_{\Lambda}/z^{n}$ does not have zeros in the same region or it vanishes identically. Lastly, it follows from (22) that the solutions $(\theta_{\Lambda}^{(n)}(z;\cdot))_{n\geq 1}$ satisfy
$\left|\theta_{\Lambda}^{(n)}(z;\boldsymbol{{x}}_{n})\right|\leq\left(C_{\beta}\,(u)\max\left\{\frac{C_{\beta}\,(u)|z|}{1-|z|/z_{0}},1\right\}\right)^{n}.$ (25)
Remark 2.2.
In fact, as will later be proved, for the Kirkwood closure process $\mathsf{K}_{z,e^{-\beta u}}$ associated to $(z,e^{-\beta u})$ there holds
$\mathsf{K}_{z,e^{-\beta u}}(N_{\Lambda}=0)=\Xi_{\Lambda}(-z)\quad\text{ for all }z\in(0,z_{0}).$
This resembles results about non-vanishing probabilities in statistical mechanics, see e.g. the fundamental theorem in [11] for the case of lattice gases.
For $z>0$ the solutions $(\theta^{(n)}_{\Lambda})_{n\geq 1}$ of the finite volume Kirkwood-Salsburg equations (20) are the correlation functions of the so-called grand canonical Gibbs measure $\mathsf{G}_{\Lambda,\beta,z,u}$ on $\Lambda$. It can be shown that when choosing a sequence of increasing sets $\Lambda_{l}\subset\Lambda_{l+1}$ such that for any bounded set $\Delta\subset\mathbb{R}^{d}$ there is an $l_{0}$ such that $\Delta\subset\Lambda_{l_{0}}$ (this limit is denoted by $\Lambda\nearrow\mathbb{R}^{d}$) these finite volume Gibbs measures converge to a limit $\mathsf{P}_{\beta,z,u}$, cf. [10]. This limit is tempered and satisfies the (multivariate) GNZ-equation (named for Georgii, Nguyen and Zessin), i.e. for every $F\colon(\mathbb{R}^{d})^{n}\times\Gamma\to[0,+\infty]$ there holds
$\int_{\Gamma}\sum_{x_{1},\dots,x_{n}\in\eta\atop x_{i}\neq x_{j}}F(\boldsymbol{{x}}_{n};\eta)\mathop{}\!\mathrm{d}\mathsf{P}_{\beta,z,u}=\int_{(\mathbb{R}^{d})^{n}}\int_{\Gamma}F(\boldsymbol{{x}}_{n};\eta\cup\{\boldsymbol{{x}}_{n}\})z^{n}e^{-\beta H(\boldsymbol{{x}}_{n})-\beta W(\{\boldsymbol{{x}}_{n}\}\mid\eta)}\mathop{}\!\mathrm{d}\mathsf{P}_{\beta,z,u}\mathop{}\!\mathrm{d}x.$ (26)
Thus, $\mathsf{P}_{\beta,z,u}$ is a so-called $(\beta,z,u)$-Gibbs measure. The function
$\kappa(\boldsymbol{{x}}_{n};\eta):=z^{n}e^{-\beta H(\boldsymbol{{x}}_{n})-\beta W(\{\boldsymbol{{x}}_{n}\}\mid\eta)}$ (27)
is also called a Papangelou kernel.
Remark 2.3.
Note that, the solution $\boldsymbol{\theta}_{\Lambda}$ of (20) also satisfies the Kirkwood-Salsburg equation without the permutation operator $\boldsymbol{\Pi}$, namely,
$(\mathbf{I}-z\boldsymbol{\chi}_{\Lambda}\boldsymbol{K})\boldsymbol{\theta}_{\Lambda}=z\boldsymbol{\chi}_{\Lambda}\boldsymbol{e}_{1}$
by construction.
3 Existence of the Kirkwood closure process
The main result can now be stated.
Theorem 3.1.
Let $\beta>0$, $0<z<z_{0}$ (with $z_{0}$ as in (21)) and $u\colon\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$ be a stable and regular pair interaction. For $\varsigma=z$ and $\phi=e^{-\beta u}$ the Kirkwood closure process $\mathsf{K}_{\varsigma,\phi}$ exists.
Remark 3.2.
It follows from the proof of Theorem 3.1 that, the correlation functions of $\mathsf{K}_{\varsigma,\phi}$ satisfy Ruelle’s bound for $\xi=ze^{\beta B}$, thus $\mathsf{K}_{\varsigma,\phi}$ is tempered. From the proof of Theorem 3.1 one also sees that the Janossy densities of $\mathsf{K}_{\varsigma,\phi}$ are given by
$j_{\Lambda}^{(n)}(z;\boldsymbol{{x}}_{n})=(-1)^{n}\Xi_{\Lambda}(-z)\theta^{(n)}_{\Lambda}(-z;\boldsymbol{{x}}_{n}).$ (28)
As previously mentioned, in computational physics the Kirkwood superposition approximation is used to approximate the correlation functions of Gibbs measures. Theorem 3.1 can be used to establish an existence result of the corresponding Kirkwood closure process under some additional decay assumptions on the pair potential.
Corollary 3.3.
Let $\beta,z>0$, $u\colon\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$ be of Lennard-Jones type, and $\mathsf{P}_{\beta,z,u}$ be a corresponding $(\beta,z,u)$-Gibbs measure with density $\rho$ and radial distribution function $g$. If $z$ is sufficiently small, the Kirkwood closure process $\mathsf{K}_{\rho,g}$ for the pair $(\rho,g)$ exists.
Proof 3.4.
It is well-known, cf. [9], that $\rho=\rho(z)$ is a decreasing function of $z$. Furthermore, for $z$ sufficiently small it was shown in [2] that there exists a Lennard-Jones type potential $v$ such that $g=e^{-v}$. Therefore, if $z$ is small enough, so is $\rho$ and the Kirkwood-closure for $(\rho,g)$ exists by Theorem 3.1.
Before proving Theorem 3.1, some auxiliary results are proved.
Proposition 3.5.
For each $n\in\mathbb{N}$ and $x_{1},\dots,x_{n}\in\Lambda^{n}$ fixed either $\theta^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n})\equiv 0$ for all $z\in B_{z_{0}}$ or $\theta^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n})/z^{n}$ does not have any zeros in $B_{z_{0}}$.
Proof 3.6.
Let $z>0$ and fix $x_{1},\dots,x_{n}\in\Lambda^{n}$.
  • 1.
    Assume first that $\theta^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n})=0$. For $n=1$ it follows from $\theta^{(1)}_{\Lambda}(z;x)\geq z/\Xi_{\Lambda}(z)$ that $z=0$ which is a contradiction. Let now $n\geq 2$, since all terms in (24) are nonnegative, this means all terms of the series have to be equal to zero, i.e.
    $\int_{\Lambda^{k}}e^{-\beta H(\{\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k}\})}\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}=0,\qquad\text{ for all }k\geq 0$
    which in particular implies $H(\{\boldsymbol{{x}}_{n}\})=+\infty$ and thus
    $e^{-\beta H(\{\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k}\})}=0,\qquad\text{ for all }k\geq 0\text{ and all }\boldsymbol{{y}}_{k}\in\Lambda^{k}.$ (29)
    Again using (24) it follows that in this case $\theta^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n})\equiv 0$ for all $z\in B_{z_{0}}$.
  • 2.
    Now assume $\theta^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n})>0$. Using (24) a straightforward calculation gives
    $\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}z}\left(\frac{\theta^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n})}{z^{n}}\right)=\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}z}\left(\frac{1}{\Xi_{\Lambda}(z)}\sum_{k=0}^{\infty}\frac{z^{k}}{k!}\int_{\Lambda^{k}}e^{-\beta H(\{\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k}\})}\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}\right)$
    $=\frac{1}{\Xi_{\Lambda}(z)}\sum_{k=0}^{\infty}\frac{z^{k}}{k!}\int_{\Lambda^{k+1}}e^{-\beta H(\{\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k+1}\})}\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k+1}-\frac{\Xi_{\Lambda}^{\prime}(z)}{\Xi_{\Lambda}(z)^{2}}\sum_{k=0}^{\infty}\frac{z^{k}}{k!}\int_{\Lambda^{k}}e^{-\beta H(\{\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k}\})}\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}.$
    It thus follows that
    $z^{n+1}\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}z}\left(\frac{\theta^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n})}{z^{n}}\right)=\int_{\Lambda}\theta^{(n+1)}_{\Lambda}(z;\boldsymbol{{x}}_{n},x)\mathop{}\!\mathrm{d}x-\theta^{(n)}_{\Lambda}(\boldsymbol{{x}}_{n})\int_{\Lambda}\theta^{(1)}_{\Lambda}(z;x)\mathop{}\!\mathrm{d}x$
    and it can be concluded that
    $\int_{\Lambda}\theta^{(n+1)}_{\Lambda}(z;\boldsymbol{{x}}_{n},x)\mathop{}\!\mathrm{d}x-\theta^{(n)}_{\Lambda}(\boldsymbol{{x}}_{n})\int_{\Lambda}\theta^{(1)}_{\Lambda}(z;x)\mathop{}\!\mathrm{d}x=\theta^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n})\,z\,\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}z}\log\left(\frac{\theta^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n})}{z^{n}}\right).$ (30)
    Since the left-hand side of (30) is analytic as product and sum of analytic functions the same follows for the right-hand side and thus the claim is proved.
Remark 3.7.
Note that $\theta^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n})=0$ for some $\boldsymbol{{x}}_{n}\in\Lambda^{n}$ also implies $\theta^{(n+k)}_{\Lambda}(z;\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})=0$ for all $k\geq 1$ and any $\boldsymbol{{y}}_{k}\in\Lambda^{k}$ by (29).
Corollary 3.8.
For any $x_{1},\dots,x_{n}\in\Lambda$ and any $z\in(0,z_{0})$ there holds
$(-1)^{n}\theta_{\Lambda}^{(n)}(-z;\boldsymbol{{x}}_{n})\geq 0.$ (31)
Proof 3.9.
By Proposition 3.5 the sign of $\theta^{(n)}_{\Lambda}(-z;\boldsymbol{{x}}_{n})$ is determined by the sign of $(-z)^{n}$. This proves the claim.
Proof of Theorem 3.1. Let $u\colon\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$ be a regular and stable pair potential and $z\in(0,z_{0})$. Let $(\rho^{(n)})_{n\geq 1}$ be defined by (5) with $\varsigma=z$ and $\phi=e^{-\beta u}$. Then the functions $(\rho^{(n)})_{n\geq 1}$ satisfy Ruelle’s bound ($\mathcal{R}_{\xi}$) with $\xi=ze^{\beta B}$. It remains to show that the inequalities (8) and (9) are satisfied for every bounded $\Lambda\subset\mathbb{R}^{d}$. For (9) this follow immediately as
$1+\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k!}\int_{\Lambda^{k}}\rho^{(k)}(\boldsymbol{{y}}_{k})\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}=\Xi_{\Lambda}(-z)$
and $\Xi_{\Lambda}(z)$ has no zeros in $B_{z_{0}}$. Finally, let $\sigma_{\Lambda}^{(0)}:=\Xi_{\Lambda}(-z)$ and for $n\geq 1$
$\sigma_{\Lambda}^{(n)}(\boldsymbol{{x}}_{n}):=(-1)^{n}\Xi_{\Lambda}(-z)\theta_{\Lambda}^{(n)}(-z;\boldsymbol{{x}}_{n})$
where $(\theta_{\Lambda}^{(n)}(-z;\cdot))_{n\geq 1}$ is the solution of (20) for $-z$. By Corollary 3.8 $\sigma_{\Lambda}^{(n)}(\boldsymbol{{x}}_{n})\geq 0$ for all $\boldsymbol{{x}}_{n}\in\Lambda^{n}$ and since
$\sigma_{\Lambda}^{(n)}(\boldsymbol{{x}}_{n})=\sum_{k=0}^{\infty}\frac{(-1)^{k}}{k!}\int_{\Lambda^{k}}\rho^{(n+k)}(\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}$
by virtue of (24), the theorem is proved. $\Box$
Remark 3.10.
Theorem 3.1 can also be extended to the case that the pair potential is not translationally invariant, i.e. $u\colon(\mathbb{R}^{d})^{2}\to\mathbb{R}\cup\{+\infty\}$. The proof works the same way, however, some additional technical assumptions on $u$ need to be made, cf. [5]
4 The Kirkwood closure process is a Gibbs point process
In this section it will be shown that the Kirkwood closure process $\mathsf{K}_{\varsigma,\phi}$ for $\varsigma=z$ and $\phi=e^{-\beta u}$ satisfies the multivariate GNZ-equation. In particular, it is shown that $\mathsf{K}_{\varsigma,\phi}$ is a Gibbs point process for the Hamiltonian
$H_{\mathsf{K}}(z;\{\boldsymbol{{x}}_{n}\}):=-\log\iota^{(n)}(z;\boldsymbol{{x}}_{n})$ (32)
with
$\iota^{(n)}(z;\boldsymbol{{x}}_{n}):=(-1)^{n}\theta^{(n)}(-z;\boldsymbol{{x}}_{n}),\qquad\boldsymbol{{x}}_{n}\in(\mathbb{R}^{d})^{n}$ (33)
and $\boldsymbol{\theta}=(\theta^{(n)})_{n\geq 1}$ the unique solution of the infinite volume Kirkwood-Salsburg equations
$(\mathbf{I}-z\boldsymbol{\Pi}\boldsymbol{K})\boldsymbol{\theta}=z\boldsymbol{e}_{1}.$ (34)
Remark 4.1.
As previously mentioned, when taking the limit $\Lambda\nearrow\mathbb{R}^{d}$ the associated solutions of (20) converge uniformly on compacts to the solution of (34), i.e. for any $n\geq 1$, $\Delta\subset\mathbb{R}^{d}$ compact there holds
$\lim_{\Lambda\nearrow\mathbb{R}^{d}}\sup_{\boldsymbol{{x}}_{n}\in\Delta^{n}}\left|\theta^{(n)}(z;\boldsymbol{{x}}_{n})-\theta^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n})\right|=0.$ (35)
This is equivalent to
$\lim_{\Lambda\nearrow\mathbb{R}^{d}}\left|\int_{(\mathbb{R}^{d})^{n}}F(\boldsymbol{{x}}_{n})\,\theta^{(n)}(z;\boldsymbol{{x}}_{n})\mathop{}\!\mathrm{d}\boldsymbol{{x}}_{n}-\int_{\Lambda^{n}}F(\boldsymbol{{x}}_{n})\,\theta^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n})\mathop{}\!\mathrm{d}\boldsymbol{{x}}_{n}\right|=0$ (36)
for any $n\geq 1$ and $F\in L^{1}((\mathbb{R}^{d})^{n})$. From (35) and Corollary 3.8 it also follows that $\iota^{(n)}\geq 0$ for all $n\geq 1$. Furthermore, (35) and Remark 2.3 imply that the solution $\boldsymbol{\theta}$ of (34) also satisfies
$(\mathbf{I}-z\boldsymbol{K})\boldsymbol{\theta}=z\boldsymbol{e}_{1}.$ (37)
In other words, there holds
$\iota^{(1)}(z;x)=z\left(1+\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k!}\int_{(\mathbb{R}^{d})^{k}}\prod_{j=1}^{k}f_{\beta}\,(x-y_{j})\iota^{(k)}(z;\boldsymbol{{y}}_{k})\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}\right)$ (38)
and
$\iota^{(n+1)}(z;x,\boldsymbol{{x}}_{n})=ze^{-\beta W(\{x\}\mid\{\boldsymbol{{x}}_{n}\})}\left(\iota^{(n)}(\boldsymbol{{x}}_{n})+\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k!}\int_{(\mathbb{R}^{d})^{k}}\prod_{j=1}^{k}f_{\beta}\,(x-y_{j})\iota^{(n+k)}(z;\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}\right).$ (39)
Furthermore, the Hamiltonian defined by (32) is stable because it follows from (25) and (35) that
$0\leq\iota^{(n)}(z;\boldsymbol{{x}}_{n})\leq\left(C_{\beta}\,(u)\max\left\{\frac{C_{\beta}\,(u)|z|}{1-|z|/z_{0}},1\right\}\right)^{n}.$ (40)
For finite configurations the interaction $W_{\mathsf{K}}$ associated to $H_{\mathsf{K}}$ is characterized by (15), i.e.
$H_{\mathsf{K}}(z;\eta\cup\gamma)=H_{\mathsf{K}}(z;\gamma)+W_{\mathsf{K}}(z;\gamma\mid\eta)+H_{\mathsf{K}}(z;\eta)$ (41)
for $\eta,\gamma\in\Gamma_{0}$. Using (41) and (32) it can be concluded that
$\frac{\iota^{(n+k)}(z;\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})}{\iota^{(k)}(z;\boldsymbol{{y}}_{k})}=e^{-H_{\mathsf{K}}(z;\boldsymbol{{x}}_{n})-W_{\mathsf{K}}(z;\{\boldsymbol{{x}}_{n}\}\mid\{\boldsymbol{{y}}_{k}\})}.$ (42)
Thus (42) can be used to define the Gibbs kernel $\kappa$ (analogous to (27)) of the Kirkwood closure process for finite configurations. However, since the Kirkwood closure process is translation invariant there holds $\mathsf{K}_{\varsigma,\phi}(N_{\mathbb{R}^{d}}(\eta)<+\infty)=0$ and thus the „typical“ $\eta$ will have infinitely many points and a way to define the interaction $W_{\mathsf{K}}$ (and thus the kernel $\kappa$) for infinite $\eta$ is needed. In light of (36), one can first look at the restriction of the Kirkwood closure process $\mathsf{K}_{\varsigma,\phi}$ to a finite volume.
Lemma 4.2.
Let $\beta>0$, $0<z<z_{0}$, $u\colon\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$ be a stable and regular pair potential, and $\mathsf{K}_{\varsigma,\phi}$ be the Kirkwood closure process for $\varsigma=z$ and $\phi=e^{-\beta u}$. Then, for any nonnegative function $F\colon(\mathbb{R}^{d})^{n}\times\Gamma\to[0,+\infty)$ and any bounded set $\Lambda\subset\mathbb{R}^{d}$ there holds
$\int_{\Gamma}\sum_{x_{1},\dots,x_{n}\in\eta_{\Lambda}\atop x_{i}\neq x_{j}}F(\boldsymbol{{x}}_{n};\eta_{\Lambda})\mathop{}\!\mathrm{d}\mathsf{K}_{\varsigma,\phi}(\eta)=\int_{\Lambda^{n}}\int_{\Gamma}F(\boldsymbol{{x}}_{n};\eta_{\Lambda}\cup\{\boldsymbol{{x}}_{n}\})\frac{j_{\Lambda}^{(N_{\Lambda}(\eta)+n)}(z;\boldsymbol{{x}}_{n},\eta_{\Lambda})}{j_{\Lambda}^{(N_{\Lambda}(\eta))}(z;\eta_{\Lambda})}\mathop{}\!\mathrm{d}\mathsf{K}_{\varsigma,\phi}(\eta)\mathop{}\!\mathrm{d}\boldsymbol{{x}}_{n}.$ (43)
Here by abuse of notation $\eta_{\Lambda}$ in the argument of $j_{\Lambda}^{(N_{\Lambda}+n)}$ (respectively $j_{\Lambda}^{(N_{\Lambda})}$) denotes the vector containing the points of $\eta_{\Lambda}$.
Proof 4.3.
Let $F\colon(\mathbb{R}^{d})^{n}\times\Gamma\to[0,+\infty)$, then by the defining property of the Janossy densities of the Kirkwood closure process there holds
$\int_{\Gamma}\sum_{x_{1},\dots,x_{n}\in\eta_{\Lambda}\atop x_{i}\neq x_{j}}F(\boldsymbol{{x}}_{n};\eta_{\Lambda})\mathop{}\!\mathrm{d}\mathsf{K}_{\varsigma,\phi}(\eta)$ $=\sum_{k=0}^{\infty}\frac{1}{k!}\int_{\Lambda^{k}}\sum_{x_{1},\dots,x_{n}\in\boldsymbol{{y}}_{k}}F(\boldsymbol{{x}}_{n};\{\boldsymbol{{y}}_{k}\})j_{\Lambda}^{(k)}(z;\boldsymbol{{y}}_{k})\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}$
$=\sum_{k=n}^{\infty}\frac{1}{k!}\int_{\Lambda^{k}}n!\sum_{1\leq i_{1}<\dots<i_{n}\leq k}F(y_{i_{1}},\dots,y_{i_{n}};\{\boldsymbol{{y}}_{k}\})j_{\Lambda}^{(k)}(z;\boldsymbol{{y}}_{k})\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}.$
Easy calculation gives
$\sum_{k=n}^{\infty}\frac{n!}{k!}\int_{\Lambda^{k}}\sum_{1\leq i_{1}<\dots<i_{n}\leq k}F(y_{i_{1}},\dots,y_{i_{n}};\{\boldsymbol{{y}}_{k}\})j_{\Lambda}^{(k)}(z;\boldsymbol{{y}}_{k})\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}$
$=\sum_{k=n}^{\infty}\frac{k(k-1)\dots(k-n+1)}{k!}\int_{\Lambda^{k}}F(\boldsymbol{{y}}_{n};\{\boldsymbol{{y}}_{k}\})j_{\Lambda}^{(k)}(z;\boldsymbol{{y}}_{k})\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}$
$=\int_{\Lambda^{n}}\sum_{k=0}^{\infty}\frac{1}{k!}\int_{\Lambda^{k}}F(\boldsymbol{{x}}_{n};\{\boldsymbol{{y}}_{k}\}\cup\{\boldsymbol{{x}}_{n}\})j_{\Lambda}^{(k+n)}(z;\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}\mathop{}\!\mathrm{d}\boldsymbol{{x}}_{n}.$
By Remark 3.7 the fraction $j_{\Lambda}^{(k+n)}(\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})/j_{\Lambda}^{(k)}(\boldsymbol{{y}}_{k})$ is well-defined and thus
$\int_{\Lambda^{n}}\sum_{k=0}^{\infty}\frac{1}{k!}\int_{\Lambda^{k}}F(\boldsymbol{{x}}_{n};\{\boldsymbol{{y}}_{k}\}\cup\{\boldsymbol{{x}}_{n}\})j_{\Lambda}^{(k+n)}(z;\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}\mathop{}\!\mathrm{d}\boldsymbol{{x}}_{n}$
=\!= $\int_{\Lambda^{n}}\sum_{k=0}^{\infty}\frac{1}{k!}\int_{\Lambda^{k}}F(\boldsymbol{{x}}_{n};\{\boldsymbol{{y}}_{k}\}\cup\{\boldsymbol{{x}}_{n}\})\frac{j_{\Lambda}^{(k+n)}(z;\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})}{j_{\Lambda}^{(k)}(z;\boldsymbol{{y}}_{k})}j_{\Lambda}^{(k)}(z;\boldsymbol{{y}}_{k})\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}\mathop{}\!\mathrm{d}\boldsymbol{{x}}_{n}.$
By the property of the Janossy densities this is equivalent to the right-hand side of (43).
Notice, that by Remark 2.3 for any bounded $\Lambda\subset\mathbb{R}^{d}$ and $z\in B_{z_{0}}$ there holds
$\frac{\theta^{(n+k)}_{\Lambda}(z;\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})}{\theta^{(k)}_{\Lambda}(z;\boldsymbol{{y}}_{k})}=ze^{-\beta W(x_{1}\mid\{\boldsymbol{{x}}_{2,n},\boldsymbol{{y}}_{k}\})}\sum_{l=0}^{\infty}\frac{1}{l!}\int_{\Lambda^{l}}\prod_{i=1}^{l}f_{\beta}(x_{1}-y_{i})\frac{\theta^{(n+k+l-1)}_{\Lambda}(z;\boldsymbol{{x}}_{2,n},\boldsymbol{{y}}_{k},{\boldsymbol{{w}}}_{l})}{\theta^{(k)}_{\Lambda}(z;\boldsymbol{{y}}_{k})}\mathop{}\!\mathrm{d}\boldsymbol{{w}}_{l}$ (44)
where $\boldsymbol{{x}}_{2,n}=(x_{2},\dots,x_{n})$ and the quotient is well-defined by Remark 3.7. From (28) it follows that the left-hand side of (44) is given by
$\frac{(-1)^{n}\theta^{(k+n)}_{\Lambda}(-z;\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})}{\theta^{(k)}_{\Lambda}(-z;\boldsymbol{{y}}_{k})}=\frac{j_{\Lambda}^{(k+n)}(z;\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})}{j_{\Lambda}^{(k)}(z;\boldsymbol{{y}}_{k})}.$ (45)
In light of Remark 4.1 one would like to take the limit $\Lambda\nearrow\mathbb{R}^{d}$ in (44) to identify the limit
$\lim_{\Lambda\nearrow\mathbb{R}^{d}}\frac{j_{\Lambda}^{(N_{\Lambda}(\eta)+n)}(z;\boldsymbol{{x}}_{n},\eta_{\Lambda})}{j_{\Lambda}^{(N_{\Lambda}(\eta))}(z;\eta_{\Lambda})}.$
As previously mentioned however, typical configurations are infinite and thus one needs to work in a function space that allows infinite configurations as arguments. Thus, the measure $\mu$ on $(\Gamma_{0}\times\Gamma,\mathscr{F}\otimes\mathscr{F})$ defined by
$\mathop{}\!\mathrm{d}\mu=\mathop{}\!\mathrm{d}\mathsf{K}_{\varsigma,\phi}+\sum_{n=1}^{\infty}\frac{C_{\beta}\,(u)^{-n}}{n!}\mathop{}\!\mathrm{d}(\lambda\!\!\!\hskip 0.5pt\lambda^{n}\times\mathsf{K}_{\varsigma,\phi})$ (46)
is introduced, as well as the spaces
$E_{\Gamma}^{1}:=L^{1}(\Gamma_{0}\times\Gamma)=\left\{\boldsymbol{F}=(F^{(n)})_{n\geq 1}\;\middle|\;F^{(n)}\colon(\mathbb{R}^{d})^{n}\times\Gamma\to\mathbb{C},\,\,\,\|\boldsymbol{F}\|_{1,\Gamma}<+\infty\right\}$
where
$\|\boldsymbol{F}\|_{1,\Gamma}:=\sum_{n=1}^{\infty}\frac{C_{\beta}\,(u)^{-n}}{n!}\int_{(\mathbb{R}^{d})^{n}}\int_{\Gamma}\left|F^{(n)}(\boldsymbol{{x}}_{n};\eta)\right|\mathop{}\!\mathrm{d}\mathsf{K}_{\varsigma,\phi}\mathop{}\!\mathrm{d}\boldsymbol{{x}}_{n}$
and
$E_{\Gamma}^{\infty}:=L^{\infty}(\Gamma_{0}\times\Gamma)=\left\{\boldsymbol{\theta}=(\theta^{(n)})_{n\geq 1}\;\middle|\;\theta^{(n)}\colon(\mathbb{R}^{d})^{n}\times\Gamma\to\mathbb{C},\,\,\,\|\boldsymbol{\theta}\|_{\infty,\Gamma}<+\infty\right\}$ (47)
where
$\|\boldsymbol{\theta}\|_{\infty,\Gamma}:=\sup_{n\geq 1}\left(C_{\beta}\,(u)^{n}\!\!\!\operatorname*{ess\,sup}_{(\boldsymbol{{x}}_{n};\eta)\in(\mathbb{R}^{d})^{n}\times\Gamma}\!|\theta^{(n)}(\boldsymbol{{x}}_{n};\eta)|\right)$
and the essential supremum is taken with respect to $\lambda\!\!\!\hskip 0.5pt\lambda^{n}\times\mathsf{K}_{\rho,g}$. In this context (44) can be rewritten as
$\kappa^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n};\eta)=ze^{-\beta W(x_{1}\mid\{\boldsymbol{{x}}_{2,n}\}\cup\eta)}\sum_{k=0}^{\infty}\frac{1}{k!}\int_{\Lambda^{k}}\prod_{i=1}^{k}f_{\beta}(x_{1}-y_{i})\kappa^{(n-1)}_{\Lambda}(z;\boldsymbol{{x}}_{2,n};\eta)\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}$ (48)
with the convention $\kappa_{\Lambda}^{(0)}\equiv 1$. Now it remains to discuss whether (48) has a solution in $E^{\infty}_{\Gamma}$ for $|z|<z_{0}$. For this two additional assumptions are made:
  • (i)
    $u\colon\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$ is locally stable, i.e. for any $n\geq 1$ and $x,x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ there holds
    $W(\{x\}\mid\{\boldsymbol{{x}}_{n}\})=\sum_{i=1}^{n}u(x_{i}-x)\geq-2B.$ (49)
  • (ii)
    $u$ is lower regular, i.e. there exists a decreasing function $\psi\colon[0,+\infty)\to[0,+\infty)$ with
    $\int_{0}^{\infty}\psi(r)r^{d-1}\mathop{}\!\mathrm{d}r<+\infty$
    and for all $x\in\mathbb{R}^{d}$
    $u(x)\geq-\psi(|x|).$
Remark 4.4.
Obviously, every locally stable stable pair potential is stable. However, neither regularity implies lower regularity nor vice versa.
Remark 4.5.
Under Assumption (ii) for any tempered point process $\mathsf{P}$ and $\eta\in\Gamma_{0}$ there holds
$W(\eta\mid\gamma)=\lim_{\Lambda\nearrow\mathbb{R}^{d}}W(\eta\mid\gamma_{\Lambda})\in\mathbb{R}\cup\{+\infty\},$ (50)
for $\mathsf{P}$-almost all $\gamma\in\Gamma,$ see [5].
Let $u\colon\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$ now be a locally stable, regular and lower regular pair interaction and define the operator $\boldsymbol{K}_{\Gamma}\colon E_{\Gamma}^{\infty}\to E_{\Gamma}^{\infty}$ by
$(\boldsymbol{K}_{\Gamma}\boldsymbol{\theta})^{(1)}(x;\eta)=e^{-\beta W(\{x\}\mid\eta)}\sum_{k=1}^{\infty}\frac{1}{k!}\int_{(\mathbb{R}^{d})^{k}}\prod_{i=1}^{k}f_{\beta}\,(x-y_{i})\theta^{(k)}(\boldsymbol{{y}}_{k};\eta)\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}$ (51)
and for $n\geq 1$ by
$\begin{split}(\boldsymbol{K}_{\Gamma}\boldsymbol{\theta})^{(n+1)}(x,\boldsymbol{{x}}_{n};\eta)=\,e^{-\beta W(\{x\}\mid\eta\cup\{\boldsymbol{{x}}_{n}\})}\left(\theta^{(n)}(\boldsymbol{{x}}_{n};\eta)+\sum_{k=1}^{\infty}\frac{1}{k!}\int_{(\mathbb{R}^{d})^{k}}\prod_{j=1}^{k}f_{\beta}\,(x-y_{j})\theta^{(n+k)}(\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k};\eta)\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}\right).\end{split}$ (52)
Then, as in Section 3, $\|\boldsymbol{K}_{\Gamma}\|_{E_{\Gamma}^{\infty}\to E_{\Gamma}^{\infty}}\leq e^{2\beta B+1}C_{\beta}\,(u)$ and thus $\boldsymbol{K}_{\Gamma}$ is well-defined.
Remark 4.6.
Note, that the assumption of local stability of $u$ is needed to ensure that the operator $\boldsymbol{K}_{\Gamma}$ is mapping into $E^{\infty}_{\Gamma}$ as the property (16) does not extend to infinite configurations.
Further, $\mathbf{I}\colon E_{\Gamma}^{\infty}\to E_{\Gamma}^{\infty}$ denotes the identity operator, and for $\Lambda\subset\mathbb{R}^{d}$ $\boldsymbol{\chi}_{\Lambda}\colon E_{\Gamma}^{\infty}\to E_{\Gamma}^{\infty}$ denotes the projection operator
$\boldsymbol{\chi}_{\Lambda}\colon$ $E_{\Gamma}^{\infty}\to E_{\Gamma}^{\infty}$
$\boldsymbol{\theta}\mapsto\boldsymbol{\chi}_{\Lambda}\boldsymbol{\theta}=(\mathds{1}_{\Lambda^{n}\times\Gamma_{\Lambda}}\theta^{(n)})_{n\geq 1}.$
Lastly, $\boldsymbol{\alpha}$ denotes the vector in $E_{\Gamma}^{\infty}$ defined by $\boldsymbol{\alpha}^{(1)}(x;\eta)=e^{-\beta W(\{x\}\mid\eta)}$ and $\boldsymbol{\alpha}^{(n)}\equiv 0$ for all $n\geq 2$. As $\|\boldsymbol{K}_{\Gamma}\|_{E_{\Gamma}^{\infty}\to E_{\Gamma}^{\infty}}\leq e^{2\beta B+1}C_{\beta}\,(u)$ the operators $\mathbf{I}-z\boldsymbol{\chi}_{\Lambda}\boldsymbol{K}_{\Gamma}$ and $\mathbf{I}-z\boldsymbol{K}_{\Gamma}$ are invertible for every $z\in B_{z_{0}}$. In particular, the equations
$(\mathbf{I}-z\boldsymbol{\chi}_{\Lambda}\boldsymbol{K}_{\Gamma})\boldsymbol{\theta}=z\boldsymbol{\chi}_{\Lambda}\boldsymbol{\alpha}$ (53)
and
$(\mathbf{I}-z\boldsymbol{K}_{\Gamma})\boldsymbol{\theta}=z\boldsymbol{\alpha}$ (54)
have unique solutions $\boldsymbol{\kappa}_{\Lambda}(z)=(\kappa_{\Lambda}^{(n)}(z;\,\cdot\,;\,\cdot\,))_{n\geq 1}$ and $\boldsymbol{\kappa}(z)=(\kappa^{(n)}(z;\,\cdot\,;\,\cdot\,))_{n\geq 1}$ in $E^{\infty}_{\Gamma}$.
Proposition 4.7.
Let $u\colon\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$ be a locally stable, regular and lower regular pair interaction and $|z|<z_{0}$. Denote by $\boldsymbol{\theta}_{\Lambda}$ the solution of (20), then for $\boldsymbol{\kappa}_{\Lambda}(z)=(\kappa_{\Lambda}^{(n)}z;\,\cdot\,;\,\cdot\,))_{n\geq 1}$ with
$\kappa_{\Lambda}^{(n)}(z;\boldsymbol{{x}}_{n};\eta):=\mathds{1}_{\Lambda^{n}\times\Gamma_{\Lambda}}(\boldsymbol{{x}}_{n};\eta)\frac{\theta^{(N(\eta)+n)}_{\Lambda}(z;\boldsymbol{{x}}_{n},\eta_{\Lambda})}{\theta^{(N(\eta))}_{\Lambda}(z;\eta_{\Lambda})}\quad\text{ for }n\geq 1$ (55)
there holds $\boldsymbol{\kappa}_{\Lambda}(z)\in E^{\infty}_{\Gamma}$. In particular, $\boldsymbol{\kappa}_{\Lambda}$ is the unique solution of (53).
Proof 4.8.
Since $\Lambda$ is bounded every $\eta\in\Gamma_{\Lambda}$ is finite. Let $\eta=\{\boldsymbol{{y}}_{k}\}$, then by
$\frac{\theta^{(n+k)}_{\Lambda}(z;\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})}{\theta^{(k)}_{\Lambda}(z;\boldsymbol{{y}}_{k})}=\frac{1}{\theta^{(k)}_{\Lambda}(z;\boldsymbol{{y}}_{k})}\frac{1}{\Xi_{\Lambda}(z)}\sum_{l=0}^{\infty}\frac{z^{n+k+l}}{l!}\int_{\Lambda^{l}}e^{-\beta H(\{\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k},{\boldsymbol{{w}}}_{l}\})}\mathop{}\!\mathrm{d}\boldsymbol{{w}}_{l}.$ (56)
Using (15) one can rewrite
$H(\{\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k},{\boldsymbol{{w}}}_{l}\})=H(\{\boldsymbol{{x}}_{n}\})+W(\{\boldsymbol{{x}}_{n}\}\mid\{\boldsymbol{{y}}_{k},{\boldsymbol{{w}}}_{l}\})+H(\{\boldsymbol{{y}}_{k},{\boldsymbol{{w}}}_{l}\}).$
Making use of (15) again and using (49) one finds
$H(\{\boldsymbol{{x}}_{n}\})+W(\{\boldsymbol{{x}}_{n}\}\mid\{\boldsymbol{{y}}_{k},{\boldsymbol{{w}}}_{l}\})=\sum_{i=1}^{n}W(\{x_{i}\}\mid\{\boldsymbol{{y}}_{k},{\boldsymbol{{w}}}_{l}\})\geq-2nB$
and it thus follows from (56) that
$\left|\frac{\theta^{(n+k)}_{\Lambda}(z;\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})}{\theta^{(k)}_{\Lambda}(z;\boldsymbol{{y}}_{k})}\right|\leq|z|^{n}e^{2\beta Bn}\left|\frac{1}{\theta^{(k)}_{\Lambda}(z;\boldsymbol{{y}}_{k})}\frac{1}{\Xi_{\Lambda}(z)}\sum_{l=0}^{\infty}\frac{z^{k+l}}{l!}\int_{\Lambda^{l}}e^{-\beta H(\{\boldsymbol{{y}}_{k},{\boldsymbol{{w}}}_{l}\})}\mathop{}\!\mathrm{d}\boldsymbol{{w}}_{l}\right|=|z|^{n}e^{2\beta Bn}.$ (57)
From (57) and (21) it follows for $\boldsymbol{\kappa}_{\Lambda}(z)$ defined by (55) that
$\|\boldsymbol{\kappa}_{\Lambda}(z)\|_{\infty,\Gamma}\leq\sup_{n\geq 1}\left(C_{\beta}\,(u)|z|e^{2\beta B}\right)^{n}<\sup_{n\geq 1}\left(\frac{1}{e}\right)^{n}<+\infty.$ (58)
Thus, $\boldsymbol{\kappa}_{\Lambda}\in E^{\infty}_{\Gamma}$ and solves (53) by (44).
From Proposition 4.7 and Remark 4.1 it also follows that for any $\eta\in\Gamma_{0}$ there holds
$\kappa^{(n)}(z;\boldsymbol{{x}}_{n};\eta)=\frac{\theta^{(N(\eta)+n)}(z;\boldsymbol{{x}}_{n};\eta)}{\theta^{(N(\eta))}(z;\eta)}\quad\text{ for }n\geq 1.$ (59)
From this it also follows that $(-1)^{n}\kappa_{\Lambda}(-z;\boldsymbol{{x}}_{n};\eta)\geq 0$ and $(-1)^{n}\kappa(-z;\boldsymbol{{x}}_{n};\eta)\geq 0$ for all $\eta\in\Gamma_{0}$, $n\geq 1$, $\boldsymbol{{x}}_{n}\in(\mathbb{R}^{d})^{n}$ and $0<z<z_{0}$.
Theorem 4.9.
Let $\beta>0,$ $0<z<z_{0}$ and $u\colon\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$ be locally stable, regular and lower regular, and let $\mathsf{K}_{\varsigma,\phi}$ be the Kirkwood closure process for $\varsigma=z$ and $\phi=e^{-\beta u}$, then for any $n\geq 1$ and any nonnegative function $F\colon(\mathbb{R})^{n}\times\Gamma\to[0,+\infty)$ there holds
$\int_{\Gamma}\sum_{x_{1},\dots,x_{n}\in\eta\atop x_{i}\neq x_{j}}F(\boldsymbol{{x}}_{n};\eta)\mathop{}\!\mathrm{d}\mathsf{K}_{\varsigma,\phi}(\eta)=\int_{(\mathbb{R}^{d})^{n}}\int_{\Gamma}F(\boldsymbol{{x}}_{n};\eta\cup\{\boldsymbol{{x}}_{n}\})(-1)^{n}\kappa^{(n)}(-z;\boldsymbol{{x}}_{n};\eta)\mathop{}\!\mathrm{d}\mathsf{K}_{\varsigma,\phi}(\eta)\mathop{}\!\mathrm{d}\boldsymbol{{x}}_{n}$ (60)
where $\kappa^{(n)}\colon(\mathbb{R}^{d})^{n}\times\Gamma\to[0,+\infty)$ is the $n$-th entry of the solution of (54). In particular, $\mathsf{K}_{\varsigma,\phi}$ satisfies the multivariate GNZ-equation and is thus a $H_{\mathsf{K}}$-Gibbs measure for $H_{\mathsf{K}}$ given by (32).
Remark 4.10.
Since the Janossy densities of the Kirkwood closure process are given by (28) it follows from (35) that for all compact $\Delta\subset\mathbb{R}^{d}$ there holds
$\lim_{\Lambda\nearrow\mathbb{R}^{d}}\sup_{\boldsymbol{{x}}_{n}\in\Delta^{n}}\left|\frac{j_{\Lambda}^{(n)}(z;\boldsymbol{{x}}_{n})}{j_{\Lambda}^{(0)}(z)}-e^{-H_{\mathsf{K}}(\{\boldsymbol{{x}}_{n}\})}\right|=0$
Proof of Theorem 4.9. Note that by (58) for any $\Lambda\subset\mathbb{R}^{d}$ there holds
$\|\boldsymbol{\kappa}_{\Lambda}(z)\|_{\Gamma_{*}}\leq\frac{1}{e}.$
One can thus find a subsequence (which is assumed to be the whole sequence) and some $\tilde{\boldsymbol{\kappa}}\in E_{\Gamma}^{\infty}$ such that
$\lim_{\Lambda\nearrow\mathbb{R}^{d}}$ $\sum_{n=1}\frac{1}{n!}\int_{(\mathbb{R}^{d})^{n}}\int_{\Gamma}F^{(n)}(\boldsymbol{{x}}_{n};\eta)\kappa^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n};\eta)\mathop{}\!\mathrm{d}\mathsf{K}_{\varsigma,\phi}\mathop{}\!\mathrm{d}\boldsymbol{{x}}_{n}$
$=$ $\sum_{n=1}\frac{1}{n!}\int_{(\mathbb{R}^{d})^{n}}\int_{\Gamma}F^{(n)}(\boldsymbol{{x}}_{n};\eta)\tilde{\kappa}^{(n)}(z;\boldsymbol{{x}}_{n};\eta)\mathop{}\!\mathrm{d}\mathsf{K}_{\varsigma,\phi}\mathop{}\!\mathrm{d}\boldsymbol{{x}}_{n}$
for each $\boldsymbol{F}=(F^{(n)})_{n\geq 1}\in E_{\Gamma}^{1}$. By (53) one finds that
$\sum_{n=1}\frac{1}{n!}$ $\int_{(\mathbb{R}^{d})^{n}}\int_{\Gamma}F^{(n)}(\boldsymbol{{x}}_{n};\eta)\kappa^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n};\eta)\mathop{}\!\mathrm{d}\mathsf{K}_{\varsigma,\phi}\mathop{}\!\mathrm{d}\boldsymbol{{x}}_{n}$
$=\sum_{n=1}\frac{1}{n!}$ $\int_{(\mathbb{R}^{d})^{n}}\int_{\Gamma}F^{(n)}(\boldsymbol{{x}}_{n};\eta)z\mathds{1}_{\Lambda^{n}\times\Gamma_{\Lambda}}(\boldsymbol{{x}}_{n};\eta)e^{-\beta W(\{x_{1}\}\mid\eta\cup\{\boldsymbol{{x}}_{2,n}\})}\cdot$
$\sum_{l=0}^{\infty}\frac{1}{l!}\int_{(\mathbb{R}^{d})^{l}}\prod_{j=1}^{l}f_{\beta}(x_{1}-y_{j})\kappa^{(n-1+l)}_{\Lambda}(z;\boldsymbol{{x}}_{2,n},\boldsymbol{{y}}_{l};\eta)\boldsymbol{{y}}_{l}\mathop{}\!\mathrm{d}\mathsf{K}_{\varsigma,\phi}\mathop{}\!\mathrm{d}\boldsymbol{{x}}_{n}$
where again $\kappa_{\Lambda}^{(0)}(\eta)\equiv 1$. Note that by Remark 4.5 there holds
$\mathds{1}_{\Lambda^{n}\times\Gamma_{\Lambda}}(\boldsymbol{{x}}_{n};\eta)e^{-\beta W(\{x_{1}\}\mid\eta\cup\{\boldsymbol{{x}}_{2,n}\})}\to e^{-\beta W(\{x_{1}\}\mid\eta\cup\{\boldsymbol{{x}}_{2,n}\})}$
pointwise $\lambda\!\!\!\hskip 0.5pt\lambda^{n}\times\mathsf{K}_{\rho,g}$-almost everywhere as $\Lambda\nearrow\mathbb{R}^{d}$, furthermore by (49) there holds $e^{-\beta W(\{x_{1}\}\mid\eta\cup\{\boldsymbol{{x}}_{2,n}\})}\leq e^{2\beta B}$ and thus by dominated convergence it can be concluded that
$\|\boldsymbol{F}\boldsymbol{\chi}_{\Lambda}\boldsymbol{\alpha}-\boldsymbol{F}\boldsymbol{\alpha}\|_{1,\Gamma}\to 0\qquad\text{ as }\Lambda\nearrow\mathbb{R}^{d}.$
It follows
$\sum_{n=1}\frac{1}{n!}$ $\int_{(\mathbb{R}^{d})^{n}}\int_{\Gamma}F^{(n)}(\boldsymbol{{x}}_{n};\eta)\tilde{\kappa}^{(n)}(z;\boldsymbol{{x}}_{n};\eta)\mathop{}\!\mathrm{d}\mathsf{K}_{\varsigma,\phi}\mathop{}\!\mathrm{d}\boldsymbol{{x}}_{n}$
$=\sum_{n=1}\frac{1}{n!}$ $\int_{(\mathbb{R}^{d})^{n}}\int_{\Gamma}F^{(n)}(\boldsymbol{{x}}_{n};\eta)ze^{-\beta W(\{x_{1}\}\mid\eta\cup\{\boldsymbol{{x}}_{2,n}\})}\cdot$
$\sum_{l=0}^{\infty}\frac{1}{l!}\int_{(\mathbb{R}^{d})^{l}}\prod_{j=1}^{l}f_{\beta}(x_{1}-y_{j})\tilde{\kappa}^{(n-1+l)}(z;\boldsymbol{{x}}_{2,n},\boldsymbol{{y}}_{l};\eta)\boldsymbol{{y}}_{l}\mathop{}\!\mathrm{d}\mathsf{K}_{\varsigma,\phi}\mathop{}\!\mathrm{d}\boldsymbol{{x}}_{n}$
for all $\boldsymbol{F}\in E_{\Gamma}^{1}$. The limit $\tilde{\boldsymbol{\kappa}}(z)$ thus satisfies (54) and since the solution of (54) is unique, one finds $\tilde{\boldsymbol{\kappa}}(z)=\boldsymbol{\kappa}(z)$. It can thus be concluded that for every $\boldsymbol{F}\in E_{\Gamma}^{1}$ there holds
$\begin{split}\lim_{\Lambda\nearrow\mathbb{R}^{d}}&\sum_{n=1}\frac{1}{n!}\int_{(\mathbb{R}^{d})^{n}}\int_{\Gamma}F^{(n)}(\boldsymbol{{x}}_{n};\eta)\kappa^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n};\eta)\mathop{}\!\mathrm{d}\mathsf{K}_{\varsigma,\phi}\mathop{}\!\mathrm{d}\boldsymbol{{x}}_{n}\\ =&\sum_{n=1}\frac{1}{n!}\int_{(\mathbb{R}^{d})^{n}}\int_{\Gamma}F^{(n)}(\boldsymbol{{x}}_{n};\eta){\kappa}^{(n)}(z;\boldsymbol{{x}}_{n};\eta)\mathop{}\!\mathrm{d}\mathsf{K}_{\varsigma,\phi}\mathop{}\!\mathrm{d}\boldsymbol{{x}}_{n}.\end{split}$ (61)
As mentioned before, by (55) and (28) one finds
$\frac{j_{\Lambda}^{(N_{\Lambda}(\eta)+n)}(z;\boldsymbol{{x}}_{n},\eta_{\Lambda})}{j_{\Lambda}^{(N_{\Lambda}(\eta))}(z;\eta_{\Lambda})}=(-1)^{n}\mathds{1}_{\Lambda^{n}\times\Gamma_{\Lambda}}(\boldsymbol{{x}}_{n};\eta)\frac{\theta^{(N(\eta)+n)}_{\Lambda}(-z;\boldsymbol{{x}}_{n};\pi(\eta))}{\theta^{(N(\eta))}_{\Lambda}(-z;\pi(\eta))}=(-1)^{n}\kappa^{(n)}_{\Lambda}(-z;\boldsymbol{{x}}_{n};\eta)$ (62)
and it follows from (43) that for $F\colon(\mathbb{R}^{d})^{n}\times\Gamma\to[0,+\infty)$ there holds
$\int_{\Gamma}\sum_{x_{1},\dots,x_{n}\in\eta_{\Lambda}\atop x_{i}\neq x_{j}}F(\boldsymbol{{x}}_{n};\eta_{\Lambda})\mathop{}\!\mathrm{d}\mathsf{K}_{\varsigma,\phi}(\eta)=\int_{(\mathbb{R}^{d})^{n}}\int_{\Gamma}\mathds{1}_{\Lambda^{n}}(\boldsymbol{{x}}_{n})F(\boldsymbol{{x}}_{n};\eta_{\Lambda})(-1)^{n}\kappa_{\Lambda}^{(n)}(-z;\boldsymbol{{x}}_{n};\eta_{\Lambda})\mathop{}\!\mathrm{d}\mathsf{K}_{\varsigma,\phi}(\eta)\mathop{}\!\mathrm{d}\boldsymbol{{x}}_{n}.$
Now (61) implies that
$\lim_{\Lambda\nearrow\mathbb{R}^{d}}$ $\int_{(\mathbb{R}^{d})^{n}}\int_{\Gamma}\mathds{1}_{\Lambda}(x)F(\boldsymbol{{x}}_{n};\eta_{\Lambda})(-1)^{n}\kappa_{\Lambda}^{(n)}(-z;\boldsymbol{{x}}_{n};\eta_{\Lambda}))\mathop{}\!\mathrm{d}\mathsf{K}_{\varsigma,\phi}(\eta)\mathop{}\!\mathrm{d}\boldsymbol{{x}}_{n}$
$=$ $\int_{(\mathbb{R}^{d})^{n}}\int_{\Gamma}F(\boldsymbol{{x}}_{n};\eta)(-1)^{n}\kappa^{(n)}(-z;\boldsymbol{{x}}_{n};\eta)\mathop{}\!\mathrm{d}\mathsf{K}_{\varsigma,\phi}(\eta)\mathop{}\!\mathrm{d}\boldsymbol{{x}}_{n}.$
On the other hand there also holds that
$\lim_{\Lambda\nearrow\mathbb{R}^{d}}\int_{\Gamma}\sum_{x_{1},\dots,x_{n}\in\eta_{\Lambda}\atop x_{i}\neq x_{j}}F(\boldsymbol{{x}}_{n};\eta_{\Lambda})\mathop{}\!\mathrm{d}\mathsf{K}_{\varsigma,\phi}(\eta)=\int_{\Gamma}\sum_{x_{1},\dots,x_{n}\in\eta\atop x_{i}\neq x_{j}}F(\boldsymbol{{x}}_{n};\eta)\mathop{}\!\mathrm{d}\mathsf{K}_{\varsigma,\phi}(\eta)$
which proves (60). $\Box$
5 Extension to higher order closures
The ansatz (5) with $\varphi=e^{-\beta u}$ can (in light of (10)) be rewritten as
$\rho^{(n)}(\boldsymbol{{x}}_{n})=\varsigma^{n}e^{-\beta H(\boldsymbol{{x}}_{n})}.$ (63)
The definition (63) continues to make sense when the Hamiltonian $H$ is not given by a simple pair interaction, but more complicated multi-body potentials, i.e. for each $n\geq 2$ there holds
$H(\boldsymbol{{x}}_{n})=\sum_{l=2}^{n}\sum_{1\leq i_{1}<\dots<i_{l}\leq n}u^{(l)}(\boldsymbol{{x}}_{i_{l}})$ (64)
for some family $(u^{(n)})_{n\geq 2}$ of $l$-body interaction potentials $u^{(l)}\colon(\mathbb{R}^{d})^{l}\to\mathbb{R}$. Here $\boldsymbol{{x}}_{i_{l}}=(x_{i_{1}},\dots,x_{i_{l}})$. The ansatz (63) with $H$ given by (64) leads to the multi-body Kirkwood-Salsburg operator, cf. [8]. The only difference to the two-body setting is the definition of the integral kernel of $\boldsymbol{K}$.
Recall, that for a Hamiltonian $H$ given by (10) the kernel of the Kirkwood-Salsburg equation is given by
$k^{(2)}(x;\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})=e^{-\beta W(\{x\}\mid\{\boldsymbol{{x}}_{n}\})}\prod_{i=1}^{k}f_{\beta}\,(x-y_{j}).$ (65)
Using (11) one can expand the product on the right-hand side of (65) to get
$k^{(2)}(x;\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})=e^{-\beta W(\{x\}\mid\{\boldsymbol{{x}}_{n}\})}\sum_{l=0}^{k}\sum_{1\leq i_{1}<\dots<i_{l}\leq k}(-1)^{k-l}\prod_{j=1}^{l}e^{-\beta u(x-y_{i_{j}})}.$
Since the interaction $W$ is linear in the second argument there holds
$W(\{x\}\mid\{\boldsymbol{{x}}_{n}\})+\sum_{j=1}^{l}u(x-y_{i_{j}})=W(\{x\}\mid\{\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{i_{l}}\})$
and thus
$k^{(2)}(x;\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})=\sum_{l=0}^{k}\sum_{1\leq i_{1}<\dots<i_{l}\leq k}(-1)^{k-l}e^{-\beta W(\{x\}\mid\{\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{i_{l}}\})}.$
Finally, it follows from (15) that this is equivalent to
$k^{(2)}(x;\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})=\sum_{l=0}^{k}\sum_{1\leq i_{1}<\dots<i_{l}\leq k}(-1)^{k-l}\exp\left(-\beta H((\{x,\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{i_{l}}\})+\beta H(\{\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{i_{l}}\})\right).$ (66)
The representation of $k^{(2)}$ via (66) continues to make sense when $H$ is given by (64), thus the kernel of the multi-obdy Kirkwood-Salsburg equations is defined as
$k^{(H)}(x;\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k}):=\sum_{l=0}^{k}\sum_{1\leq i_{1}<\dots<i_{l}\leq k}(-1)^{k-l}\exp\left(-\beta H((\{x,\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{i_{l}}\})+\beta H(\{\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{i_{l}}\})\right).$ (67)
The multi-body Kirkwood-Salsburg operator is then defined in an analogous way to (18) and (19) by
$(\boldsymbol{K}\boldsymbol{\theta})^{(1)}(x)=\sum_{k=1}^{\infty}\frac{1}{k!}\int_{(\mathbb{R}^{d})^{k}}k^{(H)}(x;\boldsymbol{{y}}_{k})\theta^{(k)}(\boldsymbol{{y}}_{k})\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}$ (68)
and for $n\geq 1$ by
$(\boldsymbol{K}\boldsymbol{\theta})^{(n+1)}(x,\boldsymbol{{x}}_{n})=k^{(H)}(x;\boldsymbol{{x}}_{n})\theta^{(n)}(\boldsymbol{{x}}_{n})+\sum_{k=1}^{\infty}\frac{1}{k!}\int_{(\mathbb{R}^{d})^{k}}k^{(H)}(x;\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})\theta^{(n+k)}(\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}.$ (69)
Now it only remains to be shown that the operator $\boldsymbol{K}$ with the kernel $k^{(H)}$ is in $\mathcal{L}(E_{\zeta})$ for some $\zeta>0$. In this case the analyticity of the solution of (20) does not depend on the particular definition of $\boldsymbol{K}$ as the solution is again given by (22) for $|z|$ sufficiently small and the proof can be repeated as in the two-body case.
Theorem 5.1.
Let $H$ be a stable Hamiltonian given by (64). If there are $\zeta,\delta>0$ such that the multi-body Kirkwood-Salsburg operator $\boldsymbol{K}\colon E_{\zeta}\to E_{\zeta}$ defined by (68) and (69) is bounded with norm $\|\boldsymbol{K}\|_{E_{\zeta}\to E_{\zeta}}\leq\delta$, then for $\varsigma<\delta$ there exists a tempered point process $\mathsf{P}$ with correlation functions $(\rho^{(n)})_{n\geq 1}$ given by (63).
Example 5.2.
The multi-body Kirkwood-Salsburg operator is bounded in the following cases:
  • Let $H$ be given by (64) with a family of $n$-body interactions $(u^{(n)})_{n\geq 2}$ where
    $u^{(2)}(x,y)=u(x-y)$
    for some stable and regular pair interaction $u\colon\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$ and nonnegative $u^{(n)}$ for $n\geq 3$ and in addition there is an $R>0$ with
    $u^{(n)}(\boldsymbol{{x}}_{n})=0,\qquad\text{ for all }n\geq 3$
    whenever there are indices $i\neq j\in\{1,\dots,n\}$ such that $|x_{i}-x_{j}|\geq R$. Then, the multi-body Kirkwood-Salsburg operator is bounded, see [13].
  • Let $H$ be given by (64) with a family of $n$-body interactions $(u^{(n)})_{n\geq 2}$ where $u^{(n)}\equiv 0$ for $n\geq 4$ and
    $u^{(2)}(x,y)=u(x-y)$
    for some stable and regular pair interaction $u\colon\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$. Concerning $u^{(3)}$ assume further, that there is a $\nu\in\mathbb{N}$ and functions $\phi_{l}\colon\mathbb{R}^{d}\to\mathbb{R}$, $1\leq l\leq\nu$, such that
    $\int_{\mathbb{R}^{d}}\left(\sum_{l=1}^{\nu}l^{2}\phi_{l}^{2}(x)\right)^{\tfrac{1}{2}}\mathop{}\!\mathrm{d}x<+\infty$
    and
    $u^{(3)}(\boldsymbol{{x}}_{3})=2\sum_{l=1}^{\nu}\phi_{l}(x_{2}-x_{1})\phi_{l}(x_{3}-x_{1}).$
    Then, the multi-body Kirkwood-Salsburg operator is bounded, cf. [12].
Remark 5.3.
When defining the operator $\boldsymbol{K}_{\Gamma}$ on an appropriate space $E_{\zeta,\Gamma}^{\infty}$ with the kernel $k^{(H)}$ one can also prove an analogous version of Theorem 4.9, provided (50) holds, e.g. the first setting of Example 5.2.

### v3

@@
 Realizability, Point processes, Gibbs point processes, Kirkwood-Salsburg equations, radial distribution function
-{AMS}
-82B21, 60G55
 1 Introduction
@@
 This question is related to an interesting inverse problem, namely, a realizability problem for point processes, see [6]:
-„Given $\rho>0$ and a nonnegative function $g$, does there exist a point process with density $\rho$ and radial distribution function $g$?“
+„Given $\rho>0$ and a nonnegative function $g$, does there exist a point process with density $\rho$ and radial distribution function gg?“
 The Kirkwood closure process is one possible ansatz for the solution of this problem.
@@
 Here $j_{\Lambda}^{(n+k)}(\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})=j_{\Lambda}^{(n+k)}(x_{1},\dots,x_{n},y_{1},\dots,y_{k})$ for brevity.
-If the point process $\mathsf{P}$ is stationary then the correlation functions are translationally invariant, and one can write $\rho^{(n)}=\rho^{n}g^{(n)}$ for appropriate functions $g^{(n)}$ depending on $n-1$ variables, where $\rho$ is the so-called intensity or density of the point process.
+If the point process $\mathsf{P}$ is stationary then the correlation functions are translationally invariant, and one can write $\rho^{(n)}=\rho^{n}g^{(n)}$ for appropriate functions $g^{(n)}$ depending on n1n-1 variables, where $\rho$ is the so-called intensity or density of the point process.
 For $n=2$ the function $g=g^{(2)}$ is the so-called radial distribution function.
@@
 where for $n=0$ the term $\rho^{(0)}=1$, see e.g. [5]. In fact, (7) can also be used to define a point process:
-Theorem A
+Theorem A.
 [Lenard [7]] Let $(\rho^{(n)})_{n\geq 1}$ be a family of nonnegative symmetric functions that satisfy ($\mathcal{R}_{\xi}$) for some $\xi>0$ such that for all $n\in\mathbb{N}$, all bounded $\Lambda\subset\mathbb{R}^{d}$ and all $\boldsymbol{{x}}_{n}\in\Lambda^{n}$
@@
 and
-$1+\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k!}\int_{\Lambda^{k}}\rho^{(k)}(\boldsymbol{{y}}_{k})\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}\geq 0$ (9)
+$1+\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k!}\int_{\Lambda^{k}}\rho^{(k)}(\boldsymbol{{y}}_{k})\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}\geq 0.$ (9)
 Then there exists a point process $\mathsf{P}$ with correlation functions $(\rho^{(n)})_{n\geq 1}$.
@@
 Defining the permutation operator $\boldsymbol{\Pi}\colon E_{C_{\beta}\,(u)}\to E_{C_{\beta}\,(u)}$ by $(\boldsymbol{\Pi}\boldsymbol{\theta})^{(n)}(\boldsymbol{{x}}_{n})=\theta^{(n)}(x_{i_{*}},\boldsymbol{{x}}_{n-1}^{\prime})$, one finds by (16) that
-$\sup_{n\geq 1}C_{\beta}\,(u)^{n}\|(\boldsymbol{\Pi}\boldsymbol{K}\boldsymbol{\theta})^{(n)}\|_{\infty}$ $\leq C_{\beta}\,(u)^{n}e^{2\beta B}\sum_{k=0}^{\infty}\frac{1}{k!}\int_{(\mathbb{R}^{d})^{k}}\prod_{j=1}^{k}|f_{\beta}\,(x_{i_{*}}-y_{j})|C_{\beta}\,(u)^{-n-k+1}\|\boldsymbol{\theta}\|\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}$
+$\sup_{n\geq 1}C_{\beta}\,(u)^{n}\|(\boldsymbol{\Pi}\boldsymbol{K}\boldsymbol{\theta})^{(n)}\|_{\infty}$ $\leq C_{\beta}\,(u)^{n}e^{2\beta B}\sum_{k=0}^{\infty}\frac{1}{k!}\int_{(\mathbb{R}^{d})^{k}}\prod_{j=1}^{k}|f_{\beta}\,(x_{i_{*}}-y_{j})|C_{\beta}\,(u)^{-n-k+1}\|\boldsymbol{\theta}\|_{C_{\beta}\,(u)}\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}$
 $\leq e^{2\beta B+1}C_{\beta}\,(u)$
@@
 $\left|\theta_{\Lambda}^{(n)}(z;\boldsymbol{{x}}_{n})\right|\leq\left(C_{\beta}\,(u)\max\left\{\frac{C_{\beta}\,(u)|z|}{1-|z|/z_{0}},1\right\}\right)^{n}.$ (25)
+Remark 2.2.
+Note that $\theta^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n})=0$ for some $\boldsymbol{{x}}_{n}\in\Lambda^{n}$ also implies $\theta^{(n+k)}_{\Lambda}(z;\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})=0$ for all $k\geq 1$ and any $\boldsymbol{{y}}_{k}\in\Lambda^{k}$ by (24) for $z\geq 0$.
 Remark 2.3.
@@
 If $z$ is sufficiently small, the Kirkwood closure process $\mathsf{K}_{\rho,g}$ for the pair $(\rho,g)$ exists.
-Proof 3.4.
+Proof 3.8.
 It is well-known, cf.
@@
 Proposition 3.9.
-For each $n\in\mathbb{N}$ and $x_{1},\dots,x_{n}\in\Lambda^{n}$ fixed either $\theta^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n})\equiv 0$ for all $z\in B_{z_{0}}$ or $\theta^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n})/z^{n}$ does not have any zeros in $B_{z_{0}}$.
-Proof 3.6.
-Let $z>0$ and fix $x_{1},\dots,x_{n}\in\Lambda^{n}$.
-  • 1.
-    Assume first that $\theta^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n})=0$.
-For $n=1$ it follows from $\theta^{(1)}_{\Lambda}(z;x)\geq z/\Xi_{\Lambda}(z)$ that $z=0$ which is a contradiction.
-Let now $n\geq 2$, since all terms in (24) are nonnegative, this means all terms of the series have to be equal to zero, i.e.
-    $\int_{\Lambda^{k}}e^{-\beta H(\{\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k}\})}\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}=0,\qquad\text{ for all }k\geq 0$
-    which in particular implies $H(\{\boldsymbol{{x}}_{n}\})=+\infty$ and thus
-    $e^{-\beta H(\{\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k}\})}=0,\qquad\text{ for all }k\geq 0\text{ and all }\boldsymbol{{y}}_{k}\in\Lambda^{k}.$ (29)
-    Again using (24) it follows that in this case $\theta^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n})\equiv 0$ for all $z\in B_{z_{0}}$.
-  • 2.
-    Now assume $\theta^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n})>0$. Using (24) a straightforward calculation gives
-    $\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}z}\left(\frac{\theta^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n})}{z^{n}}\right)=\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}z}\left(\frac{1}{\Xi_{\Lambda}(z)}\sum_{k=0}^{\infty}\frac{z^{k}}{k!}\int_{\Lambda^{k}}e^{-\beta H(\{\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k}\})}\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}\right)$
-    $=\frac{1}{\Xi_{\Lambda}(z)}\sum_{k=0}^{\infty}\frac{z^{k}}{k!}\int_{\Lambda^{k+1}}e^{-\beta H(\{\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k+1}\})}\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k+1}-\frac{\Xi_{\Lambda}^{\prime}(z)}{\Xi_{\Lambda}(z)^{2}}\sum_{k=0}^{\infty}\frac{z^{k}}{k!}\int_{\Lambda^{k}}e^{-\beta H(\{\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k}\})}\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}.$
-    It thus follows that
-    $z^{n+1}\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}z}\left(\frac{\theta^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n})}{z^{n}}\right)=\int_{\Lambda}\theta^{(n+1)}_{\Lambda}(z;\boldsymbol{{x}}_{n},x)\mathop{}\!\mathrm{d}x-\theta^{(n)}_{\Lambda}(\boldsymbol{{x}}_{n})\int_{\Lambda}\theta^{(1)}_{\Lambda}(z;x)\mathop{}\!\mathrm{d}x$
-    and it can be concluded that
-    $\int_{\Lambda}\theta^{(n+1)}_{\Lambda}(z;\boldsymbol{{x}}_{n},x)\mathop{}\!\mathrm{d}x-\theta^{(n)}_{\Lambda}(\boldsymbol{{x}}_{n})\int_{\Lambda}\theta^{(1)}_{\Lambda}(z;x)\mathop{}\!\mathrm{d}x=\theta^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n})\,z\,\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}z}\log\left(\frac{\theta^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n})}{z^{n}}\right).$ (30)
-    Since the left-hand side of (30) is analytic as product and sum of analytic functions the same follows for the right-hand side and thus the claim is proved.
-Remark 3.7.
-Note that $\theta^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n})=0$ for some $\boldsymbol{{x}}_{n}\in\Lambda^{n}$ also implies $\theta^{(n+k)}_{\Lambda}(z;\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})=0$ for all $k\geq 1$ and any $\boldsymbol{{y}}_{k}\in\Lambda^{k}$ by (29).
-Corollary 3.8.
 For any $x_{1},\dots,x_{n}\in\Lambda$ and any $z\in(0,z_{0})$ there holds
 $(-1)^{n}\theta_{\Lambda}^{(n)}(-z;\boldsymbol{{x}}_{n})\geq 0.$ (29)
-Proof 3.9.
-By Proposition 3.5 the sign of $\theta^{(n)}_{\Lambda}(-z;\boldsymbol{{x}}_{n})$ is determined by the sign of $(-z)^{n}$. This proves the claim.
+Remark 3.10.
+If $u\colon\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$ is locally stable, i.e. for any $n\geq 1$ and $x,x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ there holds
+$W(\{x\}\mid\{\boldsymbol{{x}}_{n}\})=\sum_{i=1}^{n}u(x_{i}-x)\geq-2B,$ (30)
+then Proposition 3.9 follows from the proof of Theorem 3.1 in [6].
+The idea of the proof is to approximate the potential $u$ by an appropriate potential $u_{\delta}$ that is locally stable and show that the corresponding solutions of (20) converge for $\delta\to 0$.
+Proof 3.11.
+Let $z\in(0,z_{0})$ be given choose $\delta>0$ such that $z\in(0,z_{\delta})$ where
+$z_{\delta}:=\left(e^{2\beta B+1}\exp\left(\delta/C_{\beta}\,(u)\right)C_{\beta}\,(u))\right)^{-1}$
+and define
+$u_{\delta}:=u+\infty\mathds{1}_{|x|<r_{0}}.$ (31)
+Here $r_{0}=r_{0}(\delta)>0$ is chosen such that
+$\int_{\mathbb{R}^{d}}\left|f_{\beta}\,^{\delta}(x)\right|\mathop{}\!\mathrm{d}x\leq C_{\beta}\,(u)+\delta$
+where $f_{\beta}\,^{\delta}:=e^{-\beta u_{\delta}(\cdot)}-1$.
+In particular, since $u_{\delta}\geq u$ one can use the same stability constant for $u_{\delta}$ as for $u$.
+Furthermore, if (16) holds for $u$ it also holds for $u_{\delta}$ and thus the definition of $\mathbf{\Pi}$ does not need to be changed.
+One can now define $\boldsymbol{K}_{\delta}$ as in (18) and (19) with $u_{\delta}$ in place of $u$ and gets that the corresponding version of (20) has a unique solution $\boldsymbol{\theta}_{\Lambda,\delta}$ for $|z|<z_{\delta}$ since $\|\boldsymbol{\Pi}\boldsymbol{K}_{\delta}\|_{E_{C_{\beta}\,(u)}\to E_{C_{\beta}\,(u)}}\leq e^{2\beta B+1}\exp\left(\delta/C_{\beta}\,(u)\right)C_{\beta}\,(u)$.
+From (22) it follows that
+$\|\boldsymbol{\theta}_{\Lambda,\delta}\|_{C_{\beta}\,(u)}\leq|z|C_{\beta}\,(u)\sum_{k=0}^{\infty}|z|^{k}\|\boldsymbol{\chi}_{\Lambda}\boldsymbol{\Pi}\boldsymbol{K}\|_{E_{C_{\beta}\,(u)}\to E_{C_{\beta}\,(u)}}^{k}.=\frac{zC_{\beta}\,(u)}{1-|z|e^{2\beta B+1}\exp\left(\delta/C_{\beta}\,(u)\right)C_{\beta}\,(u)}$ (32)
+and it follows that the sequence $(\boldsymbol{\theta}_{\Lambda,\delta})_{\delta>0}$ has a subsequence for $\delta\to 0$ such that for every $n\geq 0$ and $F_{n+1}\in L^{1}((\mathbb{R}^{d})^{n+1})$ there holds
+$\lim_{\delta\to 0}\int_{(\mathbb{R}^{d})^{n+1}}F_{n+1}(x,\boldsymbol{{x}}_{n})\theta_{\Lambda,\delta}^{(n+1)}(x,\boldsymbol{{x}}_{n})\mathop{}\!\mathrm{d}(x,\boldsymbol{{x}}_{n})=\int_{(\mathbb{R}^{d})^{n}}F_{n+1}(x,\boldsymbol{{x}}_{n})\theta_{\Lambda,*}^{(n+1)}(x,\boldsymbol{{x}}_{n})\mathop{}\!\mathrm{d}(x,\boldsymbol{{x}}_{n})$ (33)
+for some $\boldsymbol{\theta}_{\Lambda,*}=(\theta^{(n)}_{\Lambda,*})_{n\geq 1}$.
+Since $\boldsymbol{\theta}_{\Lambda,\delta}$ satisfies the Kirkwood-Salsburg equations for the modified potential $u_{\delta}$ one can conclude that
+$\int_{(\mathbb{R}^{d})^{n+1}}F_{n+1}(x,\boldsymbol{{x}}_{n})\theta_{\Lambda,\delta}^{(n+1)}(x,\boldsymbol{{x}}_{n})\mathop{}\!\mathrm{d}(x,\boldsymbol{{x}}_{n})$
+$=\int_{(\mathbb{R}^{d})^{n}}F_{n+1}(x,\boldsymbol{{x}}_{n})ze^{-\beta W_{\delta}(\{x\}\mid\{\boldsymbol{{x}}_{n}\})}\left(\theta_{\Lambda,\delta}^{(n)}(\boldsymbol{{x}}_{n})+\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k!}\int_{(\mathbb{R}^{d})^{k}}\prod_{j=1}^{k}f_{\beta}\,^{\delta}(x-y_{j})\theta_{\Lambda,\delta}^{(n+k)}(z;\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}\right).$
+again with the convention $\theta_{\Lambda,\delta}^{(0)}:=1$. Define
+$\widetilde{F}^{\delta}_{n+1+k}(x,\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k}):=F_{n+1}(x,\boldsymbol{{x}}_{n})ze^{-\beta W_{\delta}(\{x\}\mid\{\boldsymbol{{x}}_{n}\})}\prod_{j=1}^{k}f_{\beta}\,^{\delta}(x-y_{j})$ (34)
+then there holds
+$\lim_{\delta\to 0}\widetilde{F}^{\delta}_{n+1+k}(x,\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})=F_{n+1}(x,\boldsymbol{{x}}_{n})ze^{-\beta W(\{x\}\mid\{\boldsymbol{{x}}_{n}\})}\prod_{j=1}^{k}f_{\beta}\,^{\delta}(x-y_{j})$
+almost everywhere and furthermore by (16) there holds
+$\left|F_{n+1}(x,\boldsymbol{{x}}_{n})ze^{-\beta W_{\delta}(\{x\}\mid\{\boldsymbol{{x}}_{n}\})}\prod_{j=1}^{k}f_{\beta}\,^{\delta}(x-y_{j})\right|\leq\left|F_{n+1}(x,\boldsymbol{{x}}_{n})ze^{2\beta B}\prod_{j=1}^{k}f_{\beta}\,^{\delta}(x-y_{j})\right|$ (35)
+using dominated convergence it can thus be concluded that
+$\lim_{\delta\to 0}\int_{(\mathbb{R}^{d})^{n+1+k}}\left|\widetilde{F}^{\delta}_{n+1+k}(x,\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})-F_{n+1}(x,\boldsymbol{{x}}_{n})ze^{-\beta W(\{x\}\mid\{\boldsymbol{{x}}_{n}\})}\prod_{j=1}^{k}f_{\beta}\,^{\delta}(x-y_{j})\right|\mathop{}\!\mathrm{d}(x,\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})=0.$ (36)
+The $L^{1}$ convergence of the $\widetilde{F}^{\delta}_{n+1+k}$ together with (33) then shows that the limit $\boldsymbol{\theta}_{\Lambda,*}$ satisfies the Kirkwood-Salsburg equations for the original potential $u$ and satisfies (29) since every $\boldsymbol{\theta}_{\Lambda,\delta}$ satisfies (29) by Remark 3.10.
+Since this solution is unique the Proposition is proved.
 Proof of Theorem 3.5.
@@
 $\sigma_{\Lambda}^{(n)}(\boldsymbol{{x}}_{n}):=(-1)^{n}\Xi_{\Lambda}(-z)\theta_{\Lambda}^{(n)}(-z;\boldsymbol{{x}}_{n})$
-where $(\theta_{\Lambda}^{(n)}(-z;\cdot))_{n\geq 1}$ is the solution of (20) for $-z$.
-By Corollary 3.8 $\sigma_{\Lambda}^{(n)}(\boldsymbol{{x}}_{n})\geq 0$ for all $\boldsymbol{{x}}_{n}\in\Lambda^{n}$ and since
+where $(\theta_{\Lambda}^{(n)}(-z;\cdot))_{n\geq 1}$ is the solution of (20) for z-z.
+By Proposition 3.9 $\sigma_{\Lambda}^{(n)}(\boldsymbol{{x}}_{n})\geq 0$ for all $\boldsymbol{{x}}_{n}\in\Lambda^{n}$ and since
 $\sigma_{\Lambda}^{(n)}(\boldsymbol{{x}}_{n})=\sum_{k=0}^{\infty}\frac{(-1)^{k}}{k!}\int_{\Lambda^{k}}\rho^{(n+k)}(\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}$
@@
 for any $n\geq 1$ and $F\in L^{1}((\mathbb{R}^{d})^{n})$.
-From (35) and Corollary 3.8 it also follows that $\iota^{(n)}\geq 0$ for all $n\geq 1$.
+From (40) and Corollary LABEL:cor:pos it also follows that $\iota^{(n)}\geq 0$ for all $n\geq 1$.
 Furthermore, (40) and Remark 2.4 imply that the solution $\boldsymbol{\theta}$ of (39) also satisfies
@@
 Here by abuse of notation $\eta_{\Lambda}$ in the argument of $j_{\Lambda}^{(N_{\Lambda}+n)}$ (respectively $j_{\Lambda}^{(N_{\Lambda})}$) denotes the vector containing the points of $\eta_{\Lambda}$.
-Proof 4.3.
+Proof 4.15.
 Let $F\colon(\mathbb{R}^{d})^{n}\times\Gamma\to[0,+\infty)$, then by the defining property of the Janossy densities of the Kirkwood closure process there holds
@@
 $\int_{\Lambda^{n}}\sum_{k=0}^{\infty}\frac{1}{k!}\int_{\Lambda^{k}}F(\boldsymbol{{x}}_{n};\{\boldsymbol{{y}}_{k}\}\cup\{\boldsymbol{{x}}_{n}\})j_{\Lambda}^{(k+n)}(z;\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}\mathop{}\!\mathrm{d}\boldsymbol{{x}}_{n}$
-=\!= $\int_{\Lambda^{n}}\sum_{k=0}^{\infty}\frac{1}{k!}\int_{\Lambda^{k}}F(\boldsymbol{{x}}_{n};\{\boldsymbol{{y}}_{k}\}\cup\{\boldsymbol{{x}}_{n}\})\frac{j_{\Lambda}^{(k+n)}(z;\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})}{j_{\Lambda}^{(k)}(z;\boldsymbol{{y}}_{k})}j_{\Lambda}^{(k)}(z;\boldsymbol{{y}}_{k})\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}\mathop{}\!\mathrm{d}\boldsymbol{{x}}_{n}.$
+$=\!$ $\int_{\Lambda^{n}}\sum_{k=0}^{\infty}\frac{1}{k!}\int_{\Lambda^{k}}F(\boldsymbol{{x}}_{n};\{\boldsymbol{{y}}_{k}\}\cup\{\boldsymbol{{x}}_{n}\})\frac{j_{\Lambda}^{(k+n)}(z;\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})}{j_{\Lambda}^{(k)}(z;\boldsymbol{{y}}_{k})}j_{\Lambda}^{(k)}(z;\boldsymbol{{y}}_{k})\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}\mathop{}\!\mathrm{d}\boldsymbol{{x}}_{n}.$
 By the property of the Janossy densities this is equivalent to the right-hand side of (48).
@@
   • (i)
-    $u\colon\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$ is locally stable, i.e. for any $n\geq 1$ and $x,x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ there holds
-    $W(\{x\}\mid\{\boldsymbol{{x}}_{n}\})=\sum_{i=1}^{n}u(x_{i}-x)\geq-2B.$ (49)
+    $u\colon\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$ is locally stable.
   • (ii)
@@
 there holds $\boldsymbol{\kappa}_{\Lambda}(z)\in E^{\infty}_{\Gamma}$. In particular, $\boldsymbol{\kappa}_{\Lambda}$ is the unique solution of (57).
-Proof 4.8.
+Proof 4.20.
 Since $\Lambda$ is bounded every $\eta\in\Gamma_{\Lambda}$ is finite. Let $\eta=\{\boldsymbol{{y}}_{k}\}$, then by
@@
 $\kappa^{(n)}(z;\boldsymbol{{x}}_{n};\eta)=\frac{\theta^{(N(\eta)+n)}(z;\boldsymbol{{x}}_{n};\eta)}{\theta^{(N(\eta))}(z;\eta)}\quad\text{ for }n\geq 1.$ (63)
-From this it also follows that $(-1)^{n}\kappa_{\Lambda}(-z;\boldsymbol{{x}}_{n};\eta)\geq 0$ and $(-1)^{n}\kappa(-z;\boldsymbol{{x}}_{n};\eta)\geq 0$ for all $\eta\in\Gamma_{0}$, $n\geq 1$, $\boldsymbol{{x}}_{n}\in(\mathbb{R}^{d})^{n}$ and $0<z<z_{0}$.
+Lastly, one can conclude that $(-1)^{n}\kappa_{\Lambda}(-z;\boldsymbol{{x}}_{n};\eta)\geq 0$ and $(-1)^{n}\kappa(-z;\boldsymbol{{x}}_{n};\eta)\geq 0$ for all $\eta\in\Gamma_{0}$, $n\geq 1$, $\boldsymbol{{x}}_{n}\in(\mathbb{R}^{d})^{n}$ and $0<z<z_{0}$.
+Proof 4.21.
+Let $z\in(0,z_{0})$ and fix $x_{1},\dots,x_{n}\in\Lambda^{n}$.
+  • 1.
+    Assume first that $\theta^{(k)}_{\Lambda}(z;\boldsymbol{{y}}_{k})=0$.
+For $k=1$ it follows from $\theta^{(1)}_{\Lambda}(z;x)\geq z/\Xi_{\Lambda}(z)$ that $z=0$ which is a contradiction.
+Let now $k\geq 2$, since all terms in (24) are nonnegative, this means all terms of the series have to be equal to zero, i.e.
+    $\int_{\Lambda^{l}}e^{-\beta H(\{\boldsymbol{{y}}_{k},{\boldsymbol{{w}}}_{l}\})}\mathop{}\!\mathrm{d}\boldsymbol{{w}}_{l}=0,\qquad\text{ for all }l\geq 0$
+    which in particular implies $H(\{\boldsymbol{{y}}_{l}\})=+\infty$ and thus
+    $e^{-\beta H(\{\boldsymbol{{y}}_{k},{\boldsymbol{{w}}}_{l}\})}=0,\qquad\text{ for all }l\geq 0\text{ and all }{\boldsymbol{{w}}}_{l}\in\Lambda^{l}.$ (64)
+    Again using (24) it follows that in this case $\theta^{(k)}_{\Lambda}(z;\boldsymbol{{y}}_{k})\equiv 0$ for all $z\in B_{z_{0}}$.
+  • 2.
+    Now assume $\theta^{(k)}_{\Lambda}(z;\boldsymbol{{y}}_{k})>0$. Using (24) a straightforward calculation gives
+    $\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}z}\left(\frac{\theta^{(n)}_{\Lambda}(z;\boldsymbol{{y}}_{k})}{z^{k}}\right)=\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}z}\left(\frac{1}{\Xi_{\Lambda}(z)}\sum_{l=0}^{\infty}\frac{z^{l}}{l!}\int_{\Lambda^{l}}e^{-\beta H(\{\boldsymbol{{y}}_{k},{\boldsymbol{{w}}}_{l}\})}\mathop{}\!\mathrm{d}\boldsymbol{{w}}_{l}\right)$
+    $=\frac{1}{\Xi_{\Lambda}(z)}\sum_{l=0}^{\infty}\frac{z^{l}}{l!}\int_{\Lambda^{l+1}}e^{-\beta H(\{\boldsymbol{{y}}_{k},{\boldsymbol{{w}}}_{l+1}\})}\mathop{}\!\mathrm{d}\boldsymbol{{w}}_{l+1}-\frac{\Xi_{\Lambda}^{\prime}(z)}{\Xi_{\Lambda}(z)^{2}}\sum_{l=0}^{\infty}\frac{z^{l}}{l!}\int_{\Lambda^{l}}e^{-\beta H(\{\boldsymbol{{y}}_{k},{\boldsymbol{{w}}}_{l}\})}\mathop{}\!\mathrm{d}\boldsymbol{{w}}_{l}.$
+    It thus follows that
+    $z^{k+1}\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}z}\left(\frac{\theta^{(k)}_{\Lambda}(z;\boldsymbol{{y}}_{k})}{z^{k}}\right)=\int_{\Lambda}\theta^{(k+1)}_{\Lambda}(z;\boldsymbol{{y}}_{k})\mathop{}\!\mathrm{d}y-\theta^{(k)}_{\Lambda}(\boldsymbol{{y}}_{k})\int_{\Lambda}\theta^{(1)}_{\Lambda}(z;y)\mathop{}\!\mathrm{d}y$
+    and it can be concluded that
+    $\frac{1}{\theta^{(y)}_{\Lambda}(z;\boldsymbol{{y}}_{k})}\int_{\Lambda}\theta^{(k+1)}_{\Lambda}(z;\boldsymbol{{y}}_{k},y)\mathop{}\!\mathrm{d}y-\int_{\Lambda}\theta^{(1)}_{\Lambda}(z;y)\mathop{}\!\mathrm{d}y=\,z\,\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}z}\log\left(\frac{\theta^{(k)}_{\Lambda}(z;\boldsymbol{{y}}_{k})}{z^{k}}\right).$ (65)
+    Since the left-hand side of (65) is analytic as a solution of (57) the same follows for the right-hand side and thus the the sign of $\theta^{(k)}_{\Lambda}(z;\boldsymbol{{y}}_{k})\neq 0$ is determined by the sign of $z$.
 Theorem 4.22.